### v1

# Attention Layers Add Into Low-Dimensional Residual Subspaces

Abstract
While transformer models are widely believed to operate in high-dimensional hidden spaces, we show that attention outputs are confined to a surprisingly low-dimensional subspace, where about 60% of the directions account for 99% of the variance–a phenomenon that is induced by the attention output projection matrix and consistently observed across diverse model families and datasets. Critically, we find this low-rank structure as a fundamental cause of the prevalent dead feature problem in sparse dictionary learning, where it creates a mismatch between randomly initialized features and the intrinsic geometry of the activation space. Building on this insight, we propose a subspace-constrained training method for sparse autoencoders (SAEs), initializing feature directions into the active subspace of activations. Our approach reduces dead features from 87% to below 1% in Attention Output SAEs with 1M features, and can further extend to other sparse dictionary learning methods. Our findings provide both new insights into the geometry of attention and practical tools for improving sparse dictionary learning in large language models.
1 Introduction
Over the past years, mechanistic interpretability has shifted from a collection of proof-of-concept tools (Olsson et al., 2022; Wang et al., 2022; Meng et al., 2023; Gould et al., 2023) toward a fast-growing, scale-driven field (Templeton et al., 2024; Ameisen et al., 2025; Lindsey et al., 2025). This transformation is driven by a wave of sparse dictionary learning methods–such as sparse autoencoders (SAEs) and their variants (Cunningham et al., 2023; Bricken et al., 2023b; Lindsey et al., 2024b), transcoders (Dunefsky et al., 2024; Ge et al., 2024), and low-rank sparse attention (He et al., 2025)–that once targeted small models but are now being pushed to larger architectures and wider model families. As these approaches scale in performance and model coverage, they provide increasingly complete and fine-grained explanations of neural network behavior (Lindsey et al., 2024a; Gao et al., 2024).
However, scaling these approaches presents practical difficulties (Templeton et al., 2024; Gao et al., 2024; Mudide et al., 2025). As models and feature dictionaries grow, the number of parameters increases rapidly, driving up computational costs. At the same time, the prevalence of dead features leads to substantial waste in computation and memory (Templeton et al., 2024; Kissane et al., 2024), limiting the efficiency of interpretability methods. In this work, we identify low-rank activation structure as a major driver of dead features (Section 5.1).
In Section 4, we show that attention outputs exhibit a remarkably strong low-rank structure compared to multilayer perceptron (MLP) outputs and residual streams. Through singular value decomposition and intrinsic dimension analyses (Guth et al., 2023; Staats et al., 2025), we demonstrate that this phenomenon holds universally across layers, model families, and datasets, which is consistent with the universality hypothesis (Olah et al., 2020; Chughtai et al., 2023; Gurnee et al., 2024; Wang et al., 2025). We further trace the origin of this low-rank structure to the anisotropy of the output projection matrix $W^{O}$, which compresses the multi-head outputs into a lower-dimensional subspace.
In Section 5, we investigate how the low-rank nature of attention outputs interacts with SAE training. By evaluating the full suite of open-source SAEs from LlamaScope (He et al., 2024), we show that low intrinsic dimension strongly correlates with the number of dead features, suggesting a mismatch between random initialization and the low-dimensional geometry of the activations. To address this, we propose Active Subspace Initialization, which aligns SAE features with the active subspace of activations, substantially reducing dead features while improving reconstruction. Following Lindsey et al. (2024a) and Gao et al. (2024), we conduct scaling experiments, which further reveal that ASI achieves superior reconstruction across feature counts, and when combined with SparseAdam [^1] https://docs.pytorch.org/docs/stable/generated/torch.optim.SparseAdam.html, it achieves the best reconstruction in large scale and reduces dead features from 87% to below 1% in Attention Output SAEs with 1M features trained on Llama-3.1-8B (Dubey et al., 2024).
Furthermore, we show that Active Subspace Init can generalize to sparse replacement models (He et al., 2025; Dunefsky et al., 2024; Ameisen et al., 2025) (Section 5.4). When applied to other sparse dictionary learning methods, our initialization procedure systematically reduces the prevalence of dead parameters across architectures.
2 Related Work
2.1 Low-Rankness in Attention Mechanisms
Prior work has investigated various notions of “low-rankness” within attention mechanisms.: low-rank approximation of attention patterns (Wang et al., 2020; Tay et al., 2020; Raganato et al., 2020), low-rank parameterization for model compression (Noach and Goldberg, 2020; Hu et al., 2022), and the inherent low-rank bottleneck in single-head outputs (Bhojanapalli et al., 2020).
It is important to note that our perspective differs from these prior lines of work. We demonstrate that the multi-head self-attention outputs naturally exhibit a low-rank structure, revealing a distinct and under-explored phenomenon.
2.2 Linear Representation Hypothesis and Sparse Dictionary Learning Methods
The linear representation hypothesis posits that high-level concepts correspond to linear directions in representation space (Arora et al., 2018; Olah et al., 2020; Elhage et al., 2022; Park et al., 2024). Building on this view, a series of sparse dictionary learning methods have been proposed as interpretability tools, including sparse autoencoders and their variants (Cunningham et al., 2023; Bricken et al., 2023b; Lindsey et al., 2024b), transcoders (Dunefsky et al., 2024; Ge et al., 2024), and low-rank sparse attention (He et al., 2025). These approaches aim to decompose activations into combinations of sparsely activated features, while adopting different strategies, depending on their interpretability objectives, to predict or approximate feature activations. Importantly, this hypothesis has been shown to hold across a wide range of model scales (Templeton et al., 2024; Lieberum et al., 2024; He et al., 2024), architectures (Wang et al., 2025), and modalities (Abdulaal et al., 2024).
2.3 Dead Features in Sparse Dictionary Learning Methods
A persistent challenge in sparse dictionary learning methods is the emergence of dead features [^2] Following Bricken et al. (2023b), we define a feature as dead if it never activates over 10 million tokens in this paper. (Templeton et al., 2024; Kissane et al., 2024), which are also referred to as dead units in sparse replacement models (Dunefsky et al., 2024; Ge et al., 2024; He et al., 2025). These features contribute nothing to reconstruction quality, wasting parameters and computation. Existing approaches to mitigate this issue rely on auxiliary loss terms (Gao et al., 2024; Conerly et al., 2025) or resampling strategies (Bricken et al., 2023b) to encourage feature usage.
3 Preliminaries
3.1 Multi-Head Self-Attention and Notations
We consider a Transformer layer with multi-head self-attention (MHSA). Given input representations $X\in\mathbb{R}^{n\times d}$, where $n$ is the number of tokens and $d$ is the hidden size, each attention head $i$ computes:
$Q_{i}=XW^{Q}_{i},\quad K_{i}=XW^{K}_{i},\quad V_{i}=XW^{V}_{i},\quad W^{Q}_{i},W^{K}_{i},W^{V}_{i}\in\mathbb{R}^{d\times d_{h}},$
where $d_{h}=d/H$ is the dimensionality of each head, and $H$ is the total number of heads. The attention weights and head outputs are then given by:
$A_{i}=\mathrm{softmax}\left(\frac{Q_{i}K_{i}^{\top}}{\sqrt{d_{h}}}\right),\quad Z_{i}=A_{i}V_{i}.$
Let $Z=\text{Concat}[Z_{1},\dots,Z_{H}]\in\mathbb{R}^{n\times d}$ denote the concatenated output of all attention heads (Nanda and Bloom, 2022). The final attention output is computed by applying the output projection:
$O=ZW^{O},\quad W^{O}\in\mathbb{R}^{d\times d}.$
This formulation shows that $O$ can be viewed as the sum of low-dimensional outputs from each head, projected into the residual stream space. $O$ serves as the attention block’s contribution to the residual stream.
3.2 TopK Sparse Autoencoders
In this work, we adopt the TopK sparse autoencoder (TopK SAE) variant introduced by Gao et al. (2024). Unlike standard SAEs that impose an $\ell_{1}$ penalty, TopK SAE enforces exact sparsity by keeping only the top-$k$ activations in the latent representation for each input. Formally, given an input vector $x\in\mathbb{R}^{d}$, the encoder produces
$z=\mathrm{TopK}(W_{e}x+b_{e}),$
where $\mathrm{TopK}(v)$ sets to zero all but the largest $k$ entries of $v$. The decoder then reconstructs
$\hat{x}=W_{d}z+b_{d}.$
The model is trained to minimize the reconstruction loss, optionally augmented with an auxiliary penalty to prevent dead latents:
$\mathcal{L}_{\text{TopK-SAE}}=\|x-\hat{x}\|_{2}^{2}+\alpha\cdot\mathcal{L}_{\text{aux}},$
where $\mathcal{L}_{\text{aux}}$ is an optional term designed to penalize latents that never activate over a training period, and $\alpha$ balances reconstruction fidelity and latent utilization.
4 Low-Rank Structure of Attention Outputs
Figure 1: Across layers, model families and datasets, attention outputs exhibit dramatically lower intrinsic dimensions (details in Section 4.1) than residual streams and MLP outputs, showing that the attention layer writing into a low dimensional subspace of residual stream is a universal phenomenon.
We begin by presenting our central empirical finding: in Transformer models, attention outputs consistently display the strongest low-rank structure compared to MLP outputs and residual streams. As shown in Figure 1, attention outputs have a significantly lower intrinsic dimension. This phenomenon is remarkably robust, holding across different intermediate layers, model families and datasets. These observations highlight that the attention block modifies a subspace of the residual stream, while the MLP operates nearly on the full space.
4.1 Quantifying Low-Rankness with Relative Singular Values
We consider activation matrix $\widetilde{A}\in\mathbb{R}^{n\times d}$, where each row corresponds to one token’s activation vector, and $n$ is the number of data points, while $d$ is the dimensionality of the activation space (e.g., the hidden size of the model). Unless otherwise specified, $\widetilde{A}$ refers to the mean-centered activations. We refer readers to Appendix B for more details of activations sources.
To quantify the rank of data, we perform singular value decomposition (SVD) on $\widetilde{A}$:
$\widetilde{A}=U\Sigma V^{\top},$
[…]
To analyze the intrinsic dimension of these activations, we compute the smallest integer $k$ such that:
$\frac{\sum_{i=1}^{k}\sigma_{i}^{2}}{\sum_{i=1}^{r}\sigma_{i}^{2}}\geq\tau,$
for a given threshold $\tau\in(0,1)$. Since SVD yields the optimal low-rank approximation in terms of reconstruction error, this $k$ provides a principled way to assess how concentrated the activations are in a low-dimensional subspace [^3] We use 0.99 as the threshold in main text, results of some other thresholds are provided in Appendix D, with no influence to the conclusion. (Figure 1). We further compute the fraction of delta downstream loss recovered by different number of components. These metrics complement our central findings, offering a numerical characterization of low-rankness.
4.2 Empirical Evidence of Low-Rank Structure
Figure 2: Results for $1M$ samples from SlimPajama (Soboleva et al., 2023) fed into Llama-3.1-8B. (A) The attention outputs are the most low-rank, as indicated by the sharpest decay in singular values. (B) The fraction of loss recovered by different number of components, selected in descending order of singular values.
We draw our findings from three lines of evidence:
Low Intrinsic Dimension of Attention Outputs
Attention outputs exhibit significantly lower intrinsic dimensionality compared to other activation types, with an intrinsic dimension of around 60% of the total dimensionality. In contrast, MLP outputs and the residual streams show much higher intrinsic dimensions above 90% (Figure 1).
Rapid Spectral Decay in Attention Outputs
Attention outputs exhibit a sharply decaying singular value spectrum. This is quantitatively evidenced by the number of components retaining significant energy: only 1174 singular values exceed 5% of the maximum value in attention outputs, compared to 3276 for MLP outputs and 3943 for the residual streams (Figure 2(a)).
Efficient Downstream Loss Recovery
Attention outputs demonstrates superior dimensional efficiency in downstream task performance. Compared to zero ablation, attention outputs requires only 25.0% of dimensions to recover 95% of the downstream loss. This contrasts sharply with MLP outputs and residual streams, which require 78.1% and 85.9% of dimensions respectively to achieve the same recovery level (Figure 2(b)).
More results of these metrics across different layers, models and datasets are shown in Appendix C.
4.3 Low-Rankness of Attention Outputs Results from the Output Projection Matrix
Among all activation types, attention outputs consistently exhibits the most rapid spectral decay. To investigate whether this low-rank structure originates from the attention heads ($Z$), the output projection matrix ($W^{O}$), or their interaction, we perform a decomposition-based analysis.
Recall that the attention output is computed as $O=ZW^{O}$, where $Z\in\mathbb{R}^{n\times d}$ is the concatenated output of all attention heads, and $W^{O}\in\mathbb{R}^{d\times d}$ is a learned linear projection. To understand how the variance in $O$ is shaped, we analyze the variance of $O$ along an arbitrary unit direction $\hat{e}\in\mathbb{R}^{d}$, given by:
[…]
This expression highlights that the variance along $e$ is determined by two factors: the norm of $W^{O}e$ and the variance of $Z$ projected onto the direction $W^{O}\hat{e}$. Specifically, we can rewrite the variance as:
$\mathrm{Var}(O\hat{e})=\mathrm{Var}(Z\hat{v})\cdot\|v\|_{2}^{2}\quad\text{where}\quad v=W^{O}\hat{e},\ \hat{v}=\frac{v}{\|v\|_{2}}.$
We refer to $\mathrm{Var}(Zv)$ as the contribution of Z, capturing how much variance the head output $Z$ provides in that direction, and $\|v\|_{2}^{2}$ as the contribution of $W^{O}$, measuring how much the output projection $W^{O}$ scales or suppresses that direction.
We compute and visualize both quantities across a set of directions aligned with the right singular vectors of attention output, as shown in Figure 3. Our analysis reveals that the low-rank structure of attention outputs $O$ arises primarily from the anisotropy of $W^{O}$, which heavily compresses the output space into a lower-dimensional subspace. From a mechanistic perspective, an intuitive way to see this is that although each attention head contributes a $d_{\text{head}}$-dimensional subspace, the superposition of heads (Jermyn et al., 2024; He et al., 2025) inherently leads to overlaps among these subspaces. We note the output of the $i^{th}$ head as $\text{head}_{i}$. Consequently, the dimension of the MHSA output satisfies
[…]
$\dim\!\left(\bigcup_{i}\text{span}(\text{head}_{i})\right)\;\leq\;\sum_{i}\dim\!\big{(}\text{span}(\text{head}_{i})\big{)}\;=\;d_{\text{head}}\cdot n_{\text{head}}\;(\;=\;d_{\text{model}}\ \text{in standard MHSA)}.$
Figure 4: The number of dead features and the intrinsic dimension of each layer in Llama-3.1-8B, shows a surprising consistency: activations with lower intrinsic dimensions have more dead features.
5 Active Subspace Initialization for Sparse Autoencoders
5.1 Empirical Correlation Between Low-Rank Structure and Dead Features
To explore how low-rankness affects the interpretability of attention, we use the same framework and data as the original study to evaluate the LlamaScope SAEs (He et al., 2024) [^4] Another prominent open-source SAEs, GemmaScope (Lieberum et al., 2024), train their attention SAEs on Z rather than attention output., which provide a complete set of SAEs trained on attention output, MLP output, and residual stream. We find that the number of dead features is strongly related to intrinsic dimensions, as shown in Figure 4. This observation suggests that dead features may stem from a fundamental mismatch between the SAE’s randomly initialized weights and the geometry of the activation space.
5.2 Active Subspace Initialization for Sparse Autoencoders
To investigate this co-occurrence, we train SAEs on the attention output of Llama-3.1-8B at layer 15, initializing the SAE feature directions in the subspace spanned by the top $d_{\text{init}}$ singular vectors of the activations. As shown in Figure 5, we find that constraining the initialization to a lower-dimensional active subspace substantially decreases the number of dead features.
Based on this observation, we propose Active Subspace Initialization (ASI), a lightweight and generalizable strategy for scaling SAEs to high capacities. Let $d$ be the input dimension, $h$ the SAE hidden dimension, and $n$ the number of data points. Given activation matrices $\widetilde{A}\in\mathbb{R}^{n\times d}$, we compute the SVD:
$\widetilde{A}=U\Sigma V^{\top},\quad V\in\mathbb{R}^{d\times d}.$
We select the top $d_{\text{init}}$ right singular vectors to form the active subspace:
$V_{\text{active}}=V_{:,:d_{\text{init}}}\in\mathbb{R}^{d\times d_{\text{init}}}.$
We then initialize the SAE weights directly in this subspace:
$W_{d}\in\text{span}(V_{\text{active}}),\quad W_{e}=W_{d}^{\top}.$
Figure 5: Proportion of dead features and normalized MSE across different subspace dimensions. Random subspaces are used as the baseline, whereas only initialization with the active subspace yields improvement.
Using Active Subspace Initialization offers several benefits:
Figure 6: At a fixed number of features ($n=32768$), Active Subspace Init achieves a better reconstruction-sparsity trade-off than Base TopK and AuxK. A similar trend is observed in its impact on Language Model Downstream Loss. Note: The improvement in downstream loss is less pronounced than that in reconstruction, likely because Active Subspace Init allocates more features to the active subspace, which implicitly enhances reconstruction quality.
Enhanced Sparsity-Reconstruction Frontier Without Additional Compute
It empirically outperforms TopK on the sparsity-reconstruction frontier. It achieves slightly better results compared to the auxiliary loss approach while introducing no additional computational overhead (Figure 6).
Optimal Scaling Characteristics
Our approach demonstrates optimal scaling behavior across various model configurations. On SAEs with an expasion factor of 32 times or more, it can achieve the same or even better performance as the auxiliary loss approach with only half the number of parameters (Section 5.3).
Broad Architectural Applicability
The technique maintains applicability to diverse activation functions and architectural variants, as it operates directly on the intrinsic properties of activations. This generalizability is further explored in Section 5.4.
5.3 Scaling Laws
(a) Loss vs. Number of Parameters.
(b) Number of Dead vs. Number of Parameters.
(c) Loss vs. Number of Alive Parameters.
Figure 7: Scaling results of TopK SAEs and their variants enhanced with AuxK, Active Subspace Init, and SparseAdam–all trained on attention output from Llama-3.1-8B. (A) Loss at convergence across different feature counts: Active Subspace Init consistently achieves lower reconstruction error than TopK and AuxK. Active Subspace Init with SparseAdam achieves the best at large scale. (B) Dead features: Active Subspace Init reduces dead features compared to TopK, but still retains many at extremely large scales. Enhanced with SparseAdam, dead features can be reduced to less than 1%. (C) Loss across different number of alive features: Active Subspace Init achieves the most efficient utilization of alive features, while AuxK shows the lowest efficiency. Details in Section 5.3.
To understand how our method scales, we evaluate performance as the number of SAE features increases from 1616K to 11M, keeping other hyperparameters fixed (details in Appendix E). Figure 7 summarizes the results.
Active Subspace Init improves reconstruction.
As shown in Figure 7(a), the normalized MSE follows a smooth power-law decay with increasing feature count. Across all scales, Active Subspace Init consistently outperforms baseline TopK and AuxK, achieving superior reconstruction at equivalent feature counts.
Caveat: some dead features remain at extremely large scales in Active Subspace Init.
Figure 7(b) shows that, when scaling to extremely large feature counts, Active Subspace Init produces more dead features than AuxK. However, reconstruction performance remains better, indicating that the revived features from AuxK contribute little to actual reconstruction quality(Figure 7(c)).
Use Active Subspace Init with SparseAdam further improves performance.
Prior work (Bricken et al., 2023a) identified stale momentum as a key factor in dead feature formation. Building on this insight, we propose using SparseAdam, an optimizer specifically designed for sparse activation settings. By updating only the moments and parameters corresponding to non-zero gradients, SparseAdam naturally avoids stale momentum and thus mitigates the dead feature issue. As shown in Figures 7(a),  7(b), combining Active Subspace Init with SparseAdam substantially reduces dead features while reaching the lowest reconstruction error. While orthogonal to our initialization method, this choice provides a practical complement that further stabilizes training when scaling SAEs to very large capacities. We discuss more about stale momentum and SparseAdam in Appendix A.
5.4 Generalize to Sparse Replacement Models
Recent work by He et al. (2025) reports that Lorsa, a sparse replacement model for attention layers, exhibits a high proportion of dead parameters. We hypothesize that the low-rank structure of attention outputs contributes significantly to this phenomenon.
To test this, we apply Active Subspace Initialization to Lorsa [^5] Specifically, we initialize the feature directions (corresponding to $W^{O}$ in Lorsa) within the active subspace. Other hyperparameters are the same as He et al. (2025).. This modification reduces the proportion of dead parameters from $68.4\%$ to $40.5\%$ under identical hyperparameter settings (Table 1), while also slightly improving reconstruction error.
Table 1: Effect of Active Subspace Initialization on reducing dead parameters in attention replacement model.
Method Vanilla Active Subspace Init
Dead Parameters (%) 68.4 40.5
Normalized MSE 0.130 0.121
These results indicate that Active Subspace Initialization provides an effective strategy for mitigating dead parameters when training sparse replacement models on low-rank activations. We posit that complete elimination of dead heads may require additional mechanisms, such as the tied initialization used in SAEs to ensure alignment between feature encoding and decoding methods666”Match” means encoder can be initialized to predict relatively accurate feature activation values for decoder.. This approach has been shown to be crucial for reducing dead features in SAEs (Gao et al., 2024), and its absence in Lorsa may limit further improvement.
6 Conclusion
We identified the low-rank structure of attention outputs as a fundamental property of Transformer models and a key cause of dead features in sparse dictionary learning. Our proposed Active Subspace Initialization method addresses this by aligning SAE features with the intrinsic geometry of activations, reducing dead features while improving reconstruction quality. The approach generalizes beyond SAEs to sparse replacement models.
Appendix A Stale Momentum as Another Root Cause of Dead Features
Recent work by Bricken et al. [2023a] identifies stale momentum as a key cause to dead feature formation. Specifically, when a feature remains inactive over training steps, its associated optimizer momentum continues to accumulate. If the feature activates, the stale momentum results in disproportionately large updates, destabilizing training and potentially suppressing that feature permanently.
To directly address this, we adopt SparseAdam, an optimizer tailored for sparse activation settings, designed for more efficient use of compute and memory. SparseAdam updates both parameters and moments only when the corresponding feature is active. This could effectively prevent the harmful accumulation of stale momentum. Empirically, we observe that this change substantially reduces the rate of dead feature formation in large-scale SAE training. We believe that this is a core technique for scaling sparse dictionary methods, as stale momentum is a common problem for them.
Appendix B Activation Sources
The spectral characteristics of activations vary substantially across model architectures, datasets, and positional contexts. Below, we describe the experimental configurations used to support a broad and representative analysis.
Models
We study three large language models of different families–Llama-3.1-8B [^7] https://huggingface.co/meta-llama/Llama-3.1-8B, Qwen3-8B [^8] https://huggingface.co/Qwen/Qwen3-8B, and Gemma-2-9B [^9] https://huggingface.co/google/gemma-2-9b–all based on the Transformer architecture. This allows us to assess the robustness of spectral properties under varying model training configurations.
Datasets
To investigate how dataset diversity affects activation spectra, we select three datasets with varying linguistic and domain characteristics: (1) SlimPajama, an English corpus comprising web text, books, and other sources; (2) RedPajamaGithub, a large-scale code corpus; and (3) CCI3-Data, a Chinese dataset with broad domain coverage.
Activation Positions
Unless otherwise specified, activations are extracted from intermediate layers. For example, in LLaMA-3.1-8B (32 layers), we use activations from layer 15 (zero-indexed). We analyze four types of activations: (1) attention output, (2) MLP output, and (3) residual stream (post layer). For each type, we collect 10 million activation vectors. We empirically verify that this sample size is sufficient to produce stable and reproducible spectral analyses.
Appendix C More Low-Rank Result Across Different Models and Datasets
(a) Activation spectra for many samples from SlimPajama fed into pythia-2.8b.
(b) Activation spectra for many samples from CCI3-Data fed into Qwen3-8B.
(c) Activation spectra for many samples from RedPajamaGithub fed into Qwen3-8B.
Figure 8:
We present relative singular values and fraction of loss recovered for some other model-dataset pairs in Figure 8. Models include pythia-2.8b [^10] EleutherAI/pythia-2.8b. Datasets include RedPajamaGithub [^11] https://huggingface.co/datasets/cerebras/SlimPajama-627B and CCI3-Data [^12] https://huggingface.co/datasets/BAAI/CCI3-Data
(a) Intrinsic dimension with threshold $0.5$.
(b) Intrinsic dimension with threshold $0.9$.
(c) Intrinsic dimension with threshold $0.999$.
(d) Intrinsic dimension with threshold $0.9999$.
Figure 9: Comparison of intrinsic dimensions across different variance thresholds.
Appendix D Different Choose of Variance Threshold for Intrinsic Dimension
We use 0.99 as the variance threshold in the main text. We show other threshold chose make no influence to the conclusion in Figure 9. Attention outputs show low-rank structure consistently.
Appendix E SAE Training Details
E.1 Collecting Activations
We truncate each document to 1024 tokens and prepend a <bos> token to the beginning of each document. During training, we exclude the activations corresponding to the <bos> and <eos> tokens.
It has been observed that activations from different sequence positions within the same document are often highly correlated and may lack diversity. To mitigate this issue, it is common to introduce randomness into the training data. Our shuffling strategy maintains a buffer that is reshuffled whenever the buffer is refilled.
E.2 Initialization and Optimization
The decoder columns $W^{dec}_{:,i}$ are initialized uniformly and normalized to achieve the lowest initial reconstruction loss. We find that the specific initialization norm has little impact, as long as in a reasonable scope. For example, initializing $W^{dec}_{:,i}$ uniformly with a fixed bound, as in Conerly et al. [2025], yields similar results. The encoder weights $W^{enc}$ are initialized as the transpose of $W^{dec}$, while both the encoder bias $b^{enc}$ and decoder bias $b^{dec}$ are set to zero.
This initialization scheme ensures that the SAE begins training with an almost zero reconstruction loss. Such initialization has been widely observed to benefit SAE training.
We train SAEs using the Adam and SparseAdam optimizers, both with $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon=10^{-8}$.
E.3 Fixed Hyperparameters in Scaling Law
Model, Dataset, Layer, Pos
Llama-3.1-8B, SlimPajama, 15(index start at 0), attention output.
Sparsity
We empirically set $k=50$ for a reasonable sparsity in scaling laws.
Batch Size
We empirically set $batchsize=4096$, which belows the critical batch size.
Learning Rate
The learning rate for Adam and SparseAdam is sweeped separately in $1\mathrm{e}{-5}$, $2\mathrm{e}{-5}$, $4\mathrm{e}{-5}$, $6\mathrm{e}{-5}$, $8\mathrm{e}{-5}$, $1\mathrm{e}{-4}$, $2\mathrm{e}{-4}$, $4\mathrm{e}{-4}$], and we ultimately use $4\mathrm{e}{-5}$ for Adam and $6\mathrm{e}{-5}$ for SparseAdam.
AuxK
We follow Gao et al. [2024] to set auxiliary loss coefficient $\alpha$ as $\frac{1}{32}$. We sweep the $k_{aux}$ in [256, 512, 1024, 2048] and finally choose 512.
Dimension of Subspace for SAE Initialization
As shown in Figure 5, $d_{init}$ is a hyperparameter with a wide range of sub-optimal value space (from 256 to 2048). We use 768 for all experiments.
Total Tokens
We use $2.5B$ tokens for each SAE training.

### v2

@@
 Abstract
-While transformer models are widely believed to operate in high-dimensional hidden spaces, we show that attention outputs are confined to a surprisingly low-dimensional subspace, where about 60% of the directions account for 99% of the variance–a phenomenon that is induced by the attention output projection matrix and consistently observed across diverse model families and datasets.
-Critically, we find this low-rank structure as a fundamental cause of the prevalent dead feature problem in sparse dictionary learning, where it creates a mismatch between randomly initialized features and the intrinsic geometry of the activation space.
+Transformer architectures, and their attention mechanisms in particular, form the foundation of modern large language models.
+While transformer models are widely believed to operate in high-dimensional hidden spaces, we show that attention outputs are confined to a surprisingly low-dimensional subspace, where about 60% of the directions account for 99% of the variance–a phenomenon that is consistently observed across diverse model families and datasets, and is induced by the attention output projection matrix.
+Critically, we find this low-rank structure as a key factor of the prevalent dead feature problem in sparse dictionary learning, where it creates a mismatch between randomly initialized features and the intrinsic geometry of the activation space.
 Building on this insight, we propose a subspace-constrained training method for sparse autoencoders (SAEs), initializing feature directions into the active subspace of activations.
@@
 1 Introduction
-Over the past years, mechanistic interpretability has shifted from a collection of proof-of-concept tools (Olsson et al., 2022; Wang et al., 2022; Meng et al., 2023; Gould et al., 2023) toward a fast-growing, scale-driven field (Templeton et al., 2024; Ameisen et al., 2025; Lindsey et al., 2025).
-This transformation is driven by a wave of sparse dictionary learning methods–such as sparse autoencoders (SAEs) and their variants (Cunningham et al., 2023; Bricken et al., 2023b; Lindsey et al., 2024b), transcoders (Dunefsky et al., 2024; Ge et al., 2024), and low-rank sparse attention (He et al., 2025)–that once targeted small models but are now being pushed to larger architectures and wider model families.
+Over the past years, mechanistic interpretability has shifted from a collection of proof-of-concept tools (Olsson et al., 2022; Wang et al., 2022; Meng et al., 2023; Gould et al., 2023) toward a fast-growing, scale-driven field (Ameisen et al., 2025; Lindsey et al., 2025).
+This transformation is driven by a wave of sparse dictionary learning methods–such as sparse autoencoders (SAEs) and their variants (Cunningham et al., 2023; Bricken et al., 2023b; Lindsey et al., 2024b), transcoders (Dunefsky et al., 2024; Ge et al., 2024), and low-rank sparse attention (He et al., 2025)–that once targeted small models but are now being pushed to larger architectures and wider model families (Templeton et al., 2024; Gao et al., 2024; Hazra et al., 2025).
 As these approaches scale in performance and model coverage, they provide increasingly complete and fine-grained explanations of neural network behavior (Lindsey et al., 2024a; Gao et al., 2024).
@@
 In Section 4, we show that attention outputs exhibit a remarkably strong low-rank structure compared to multilayer perceptron (MLP) outputs and residual streams.
-Through singular value decomposition and intrinsic dimension analyses (Guth et al., 2023; Staats et al., 2025), we demonstrate that this phenomenon holds universally across layers, model families, and datasets, which is consistent with the universality hypothesis (Olah et al., 2020; Chughtai et al., 2023; Gurnee et al., 2024; Wang et al., 2025).
+Through singular value decomposition and intrinsic dimensionality analyses (Guth et al., 2023; Staats et al., 2025), we demonstrate that this phenomenon holds universally across layers, datasets, and model families—Llama 3.1 (Dubey et al., 2024), Gemma 2 (Rivière et al., 2024), and Qwen 3 (Yang et al., 2025), which is consistent with the universality hypothesis (Olah et al., 2020; Chughtai et al., 2023; Gurnee et al., 2024; Wang et al., 2025).
 We further trace the origin of this low-rank structure to the anisotropy of the output projection matrix $W^{O}$, which compresses the multi-head outputs into a lower-dimensional subspace.
 In Section 5, we investigate how the low-rank nature of attention outputs interacts with SAE training.
-By evaluating the full suite of open-source SAEs from LlamaScope (He et al., 2024), we show that low intrinsic dimension strongly correlates with the number of dead features, suggesting a mismatch between random initialization and the low-dimensional geometry of the activations.
-To address this, we propose Active Subspace Initialization, which aligns SAE features with the active subspace of activations, substantially reducing dead features while improving reconstruction.
+By evaluating the full suite of open-source SAEs from LlamaScope (He et al., 2024), we show that low intrinsic dimensionality strongly correlates with the number of dead features, suggesting a mismatch between random initialization and the low-dimensional geometry of the activations.
+Drawing inspiration from Phan et al.
+(2025)’s principal component initialization for the first network layer, we propose Active Subspace Initialization, which aligns SAE features with the active subspace of activations, substantially reducing dead features while improving reconstruction.
 Following Lindsey et al.
@@
 2.1 Low-Rankness in Attention Mechanisms
-Prior work has investigated various notions of “low-rankness” within attention mechanisms.: low-rank approximation of attention patterns (Wang et al., 2020; Tay et al., 2020; Raganato et al., 2020), low-rank parameterization for model compression (Noach and Goldberg, 2020; Hu et al., 2022), and the inherent low-rank bottleneck in single-head outputs (Bhojanapalli et al., 2020).
-It is important to note that our perspective differs from these prior lines of work.
-We demonstrate that the multi-head self-attention outputs naturally exhibit a low-rank structure, revealing a distinct and under-explored phenomenon.
+Prior work has investigated various notions of “low-rankness” within attention mechanisms.: low-rank approximation of attention patterns (Wang et al., 2020; Tay et al., 2020; Raganato et al., 2020), low-rank parameterization for model compression (Noach & Goldberg, 2020; Hu et al., 2022), and the inherent low-rank bottleneck in single-head outputs (Bhojanapalli et al., 2020).
+Different from these prior lines of work, we demonstrate that the multi-head self-attention outputs exhibit a low-rank structure, revealing a distinct and under-explored phenomenon.
 2.2 Linear Representation Hypothesis and Sparse Dictionary Learning Methods
-The linear representation hypothesis posits that high-level concepts correspond to linear directions in representation space (Arora et al., 2018; Olah et al., 2020; Elhage et al., 2022; Park et al., 2024).
-Building on this view, a series of sparse dictionary learning methods have been proposed as interpretability tools, including sparse autoencoders and their variants (Cunningham et al., 2023; Bricken et al., 2023b; Lindsey et al., 2024b), transcoders (Dunefsky et al., 2024; Ge et al., 2024), and low-rank sparse attention (He et al., 2025).
-These approaches aim to decompose activations into combinations of sparsely activated features, while adopting different strategies, depending on their interpretability objectives, to predict or approximate feature activations.
-Importantly, this hypothesis has been shown to hold across a wide range of model scales (Templeton et al., 2024; Lieberum et al., 2024; He et al., 2024), architectures (Wang et al., 2025), and modalities (Abdulaal et al., 2024).
+The linear representation hypothesis suggests that high-level concepts align with linear directions in representation space (Arora et al., 2018; Olah et al., 2020; Elhage et al., 2022; Park et al., 2024).
+Motivated by this view, various sparse dictionary learning methods have been developed for interpretability, including sparse autoencoders and their variants (Cunningham et al., 2023; Bricken et al., 2023b; Lindsey et al., 2024b), transcoders (Dunefsky et al., 2024; Ge et al., 2024), and low-rank sparse attention (He et al., 2025).
+These methods decompose activations into sparse feature combinations, while differing in their strategies for predicting or approximating feature activations.
+The hypothesis has been validated across diverse model scales (Templeton et al., 2024; Lieberum et al., 2024; He et al., 2024), architectures (Wang et al., 2025), and modalities (Abdulaal et al., 2024).
 2.3 Dead Features in Sparse Dictionary Learning Methods
@@
 Existing approaches to mitigate this issue rely on auxiliary loss terms (Gao et al., 2024; Conerly et al., 2025) or resampling strategies (Bricken et al., 2023b) to encourage feature usage.
+2.4 PCA-Inspired Network Initialization
+A common practice applies PCA to input data for dimensionality reduction before network training (Hastie et al., 2009; Montavon et al., 2012; Jolliffe, 1986; Bishop & Nasrabadi, 2007).
+Recently, Phan et al.
+(2025) proposed PCsInit, which initializes the first layer weights of networks with top principal components of data—embedding the PCA transform directly into the network.
+This provides the model with a superior parameter set (Gu et al., 2025), boosting performance by construction.
 3 Preliminaries
 3.1 Multi-Head Self-Attention and Notations
-We consider a Transformer layer with multi-head self-attention (MHSA).
-Given input representations $X\in\mathbb{R}^{n\times d}$, where $n$ is the number of tokens and $d$ is the hidden size, each attention head $i$ computes:
+We consider a Transformer block with multi-head self-attention (MHSA) (Vaswani et al., 2017).
+Given input activations $X\in\mathbb{R}^{n\times d}$, where $n$ is the token count and $d$ is the model hidden size, each attention head $i$ computes:
 $Q_{i}=XW^{Q}_{i},\quad K_{i}=XW^{K}_{i},\quad V_{i}=XW^{V}_{i},\quad W^{Q}_{i},W^{K}_{i},W^{V}_{i}\in\mathbb{R}^{d\times d_{h}},$
@@
 $A_{i}=\mathrm{softmax}\left(\frac{Q_{i}K_{i}^{\top}}{\sqrt{d_{h}}}\right),\quad Z_{i}=A_{i}V_{i}.$
-Let $Z=\text{Concat}[Z_{1},\dots,Z_{H}]\in\mathbb{R}^{n\times d}$ denote the concatenated output of all attention heads (Nanda and Bloom, 2022).
+Let $Z=\text{Concat}[Z_{1},\dots,Z_{H}]\in\mathbb{R}^{n\times d}$ denote the concatenated output of all attention heads (Nanda & Bloom, 2022).
 The final attention output is computed by applying the output projection:
@@
 This formulation shows that $O$ can be viewed as the sum of low-dimensional outputs from each head, projected into the residual stream space.
-$O$ serves as the attention block’s contribution to the residual stream.
+$O$ is the attention block’s contribution to the residual stream.
 3.2 TopK Sparse Autoencoders
-In this work, we adopt the TopK sparse autoencoder (TopK SAE) variant introduced by Gao et al.
+In this work, we adopt the TopK sparse autoencoder (TopK SAE) introduced by Gao et al.
 (2024).
@@
 Unless otherwise specified, $\widetilde{A}$ refers to the mean-centered activations.
-We refer readers to Appendix B for more details of activations sources.
+We refer readers to Appendix D for more details of activations sources.
 To quantify the rank of data, we perform singular value decomposition (SVD) on $\widetilde{A}$:
@@
 for a given threshold $\tau\in(0,1)$.
-Since SVD yields the optimal low-rank approximation in terms of reconstruction error, this $k$ provides a principled way to assess how concentrated the activations are in a low-dimensional subspace [^3] We use 0.99 as the threshold in main text, results of some other thresholds are provided in Appendix D, with no influence to the conclusion.
+Since SVD yields the optimal low-rank approximation in terms of reconstruction error, this $k$ provides a principled way to assess how concentrated the activations are in a low-dimensional subspace [^3] We use 0.99 as the threshold in main text, results of some other thresholds are provided in Appendix F, with no influence to the conclusion.
 (Figure 1).
-We further compute the fraction of delta downstream loss recovered by different number of components.
+We further compute the fraction of delta downstream loss recovered by different number of components (Section 4.2).
 These metrics complement our central findings, offering a numerical characterization of low-rankness.
+For each type of activations, we collect 10 million activation vectors [^4] In rare cases, outlier activations inflate variance along some directions, making variance-based dimensionality estimates unreliable; to address this, we filter activations with a norm exceeding $5\sigma$ from the mean..
+We empirically verify that this sample size is sufficient to produce stable and reproducible singular spectrum analysis.
 4.2 Empirical Evidence of Low-Rank Structure
-Figure 2: Results for $1M$ samples from SlimPajama (Soboleva et al., 2023) fed into Llama-3.1-8B.
+Figure 2: Results for 1010M tokens from SlimPajama (Soboleva et al., 2023) fed into Llama-3.1-8B.
 (a) The attention outputs are the most low-rank, as indicated by the sharpest decay in singular values.
-(B) The fraction of loss recovered by different number of components, selected in descending order of singular values.
+(b) Fraction of loss recovered using varying numbers of top singular components.
 We draw our findings from three lines of evidence:
 Low Intrinsic Dimension of Attention Outputs
-Attention outputs exhibit significantly lower intrinsic dimensionality compared to other activation types, with an intrinsic dimension of around 60% of the total dimensionality.
-In contrast, MLP outputs and the residual streams show much higher intrinsic dimensions above 90% (Figure 1).
-Rapid Spectral Decay in Attention Outputs
-Attention outputs exhibit a sharply decaying singular value spectrum.
-This is quantitatively evidenced by the number of components retaining significant energy: only 1174 singular values exceed 5% of the maximum value in attention outputs, compared to 3276 for MLP outputs and 3943 for the residual streams (Figure 2(a)).
+Attention outputs has an intrinsic dimension of around 60% of the total dimensionality. In contrast, MLP outputs and the residual streams show much higher intrinsic dimensions above 90% (Figure 1).
+Rapid Singular Spectral Decay in Attention Outputs
+This is quantitatively evidenced by the number of components retaining significant energy: only 74.7% singular values exceed 1% of the maximum in attention outputs, versus 100.0% for MLP outputs and residual streams (Figure 2(a)).
 Efficient Downstream Loss Recovery
-Attention outputs demonstrates superior dimensional efficiency in downstream task performance.
-Compared to zero ablation, attention outputs requires only 25.0% of dimensions to recover 95% of the downstream loss.
-This contrasts sharply with MLP outputs and residual streams, which require 78.1% and 85.9% of dimensions respectively to achieve the same recovery level (Figure 2(b)).
-More results of these metrics across different layers, models and datasets are shown in Appendix C.
+Compared to zero ablation, attention outputs requires only 39.1% of dimensions to recover over 99% of the downstream loss, versus 95.3% and 96.9% for MLP outputs and residual streams (Figure 2(b)).
+More results of these metrics across different layers, models and datasets are shown in Appendix E.
 4.3 Low-Rankness of Attention Outputs Results from the Output Projection Matrix
@@
 Recall that the attention output is computed as $O=ZW^{O}$, where $Z\in\mathbb{R}^{n\times d}$ is the concatenated output of all attention heads, and $W^{O}\in\mathbb{R}^{d\times d}$ is a learned linear projection.
-To understand how the variance in $O$ is shaped, we analyze the variance of $O$ along an arbitrary unit direction $\hat{e}\in\mathbb{R}^{d}$, given by:
+To understand how the variance in $O$ (singular value spectra of O) is shaped, we analyze the variance of $O$ along an arbitrary unit direction $\hat{e}\in\mathbb{R}^{d}$, given by:
 […]
@@
 $\dim\!\left(\bigcup_{i}\text{span}(\text{head}_{i})\right)\;\leq\;\sum_{i}\dim\!\big(\text{span}(\text{head}_{i})\big)\;=\;d_{\text{head}}\cdot n_{\text{head}}\;(\;=\;d_{\text{model}}\ \text{in standard MHSA)}.$
-Figure 4: The number of dead features and the intrinsic dimension of each layer in Llama-3.1-8B, shows a surprising consistency: activations with lower intrinsic dimensions have more dead features.
 5 Active Subspace Initialization for Sparse Autoencoders
 5.1 Empirical Correlation Between Low-Rank Structure and Dead Features
-To explore how low-rankness affects the interpretability of attention, we use the same framework and data as the original study to evaluate the LlamaScope SAEs (He et al., 2024) [^4] Another prominent open-source SAEs, GemmaScope (Lieberum et al., 2024), train their attention SAEs on Z rather than attention output., which provide a complete set of SAEs trained on attention output, MLP output, and residual stream.
+Figure 4: The number of dead features (left) and the intrinsic dimension (right) of each activation in Llama-3.1-8B, shows a surprising consistency: activations with lower intrinsic dimensions have more dead features, corresponding to all layers of attention output and last two layers of MLP output.
+To explore how low-rankness affects the interpretability of attention, we use the same framework and data as the original study to evaluate the LlamaScope SAEs (He et al., 2024), which provide a complete set of SAEs trained on attention output, MLP output, and residual stream  [^5] Another prominent open-source SAEs, GemmaScope (Lieberum et al., 2024), train their attention SAEs on Z rather than attention output..
 We find that the number of dead features is strongly related to intrinsic dimensions, as shown in Figure 4.
-This observation suggests that dead features may stem from a fundamental mismatch between the SAE’s randomly initialized weights and the geometry of the activation space.
+This observation suggests that dead features may stem from the low-rank geometry of the activation space.
 5.2 Active Subspace Initialization for Sparse Autoencoders
-To investigate this co-occurrence, we train SAEs on the attention output of Llama-3.1-8B at layer 15, initializing the SAE feature directions in the subspace spanned by the top $d_{\text{init}}$ singular vectors of the activations.
-As shown in Figure 5, we find that constraining the initialization to a lower-dimensional active subspace substantially decreases the number of dead features.
 Based on this observation, we propose Active Subspace Initialization (ASI), a lightweight and generalizable strategy for scaling SAEs to high capacities.
-Let $d$ be the input dimension, $h$ the SAE hidden dimension, and $n$ the number of data points.
-Given activation matrices $\widetilde{A}\in\mathbb{R}^{n\times d}$, we compute the SVD:
-$\widetilde{A}=U\Sigma V^{\top},\quad V\in\mathbb{R}^{d\times d}.$
-We select the top $d_{\text{init}}$ right singular vectors to form the active subspace:
-$V_{\text{active}}=V_{:,:d_{\text{init}}}\in\mathbb{R}^{d\times d_{\text{init}}}.$
-We then initialize the SAE weights directly in this subspace:
-$W_{d}\in\text{span}(V_{\text{active}}),\quad W_{e}=W_{d}^{\top}.$
-Figure 5: Proportion of dead features and normalized MSE across different subspace dimensions.
-Random subspaces are used as the baseline, whereas only initialization with the active subspace yields improvement.
+Let $d$ denote the input dimension, $h$ the hidden dimension of the SAE, and $n$ the number of data points.
+Given activation matrices $\widetilde{A}\in\mathbb{R}^{n\times d}$, we compute the singular value decomposition (SVD):
+$\widetilde{A}=U\Sigma V^{\top},\quad V\in\mathbb{R}^{d\times d},$
+where $U\in\mathbb{R}^{n\times d}$ is the left singular vector matrix, $\Sigma\in\mathbb{R}^{d\times d}$ is the diagonal matrix of singular values, and $V\in\mathbb{R}^{d\times d}$ contains the right singular vectors.
+We select the top $d_{\text{init}}$ right singular vectors to form the active subspace and then initialize the SAE weights directly in this subspace:
+$V_{\text{active}}=V_{:,:d_{\text{init}}}\in\mathbb{R}^{d\times d_{\text{init}}},\quad W_{D}\in\text{span}(V_{\text{active}}),\quad W_{E}=W_{D}^{\top},$
+where $W_{E}$ is the encoder weight matrix and $W_{D}$ is the decoder weight matrix.
+Intuitively, ASI aligns the initial SAE parameters with the active directions of the data, ensuring that SAEs start in a meaningful low-dimensional subspace.
+As $d_{\text{init}}$ decreases from the full space dimension [^6] Setting $D_{init}$ equal to the full space dimension is equivalent to not using Active Subspace Initialization.
+within a certain range, the number of dead features in the SAE rapidly drops, with a corresponding improvement in Delta LM loss [^7] This metric is defined as the difference between the original language model loss and the loss when the SAE is inserted at the corresponding position, evaluated over 1 million tokens.
+(Figure 5).
+Further ablation studies and implementation details are contained in Appendix B and Appendix I, respectively.
+Figure 5: After using ASI, proportion of dead features (left), normalized MSE (mid) and Delta LM loss (right) across different subspace dimensions for activations with a full space dimension of 4096.
+Figure 6: At a fixed number of features ($n=32768$), Active Subspace Init achieves a better reconstruction-sparsity trade-off than TopK and AuxK.
+A similar trend is observed in its impact on Delta LM Loss.
+Note: The improvement in downstream loss is less pronounced than that in reconstruction, likely because Active Subspace Init allocates more features to the active subspace, which implicitly enhances reconstruction quality.
 Using Active Subspace Initialization offers several benefits:
-Figure 6: At a fixed number of features ($n=32768$), Active Subspace Init achieves a better reconstruction-sparsity trade-off than Base TopK and AuxK.
-A similar trend is observed in its impact on Language Model Downstream Loss.
-Note: The improvement in downstream loss is less pronounced than that in reconstruction, likely because Active Subspace Init allocates more features to the active subspace, which implicitly enhances reconstruction quality.
-Enhanced Sparsity-Reconstruction Frontier Without Additional Compute
-It empirically outperforms TopK on the sparsity-reconstruction frontier.
-It achieves slightly better results compared to the auxiliary loss approach while introducing no additional computational overhead (Figure 6).
+Reduced dead features and Enhanced Sparsity-Reconstruction Frontier Without Additional Compute
+It achieves near-zero dead features and slightly superior results compared to the auxiliary loss approach (AuxK), at no additional computational cost of the same order. (Figure 6).
 Optimal Scaling Characteristics
-Our approach demonstrates optimal scaling behavior across various model configurations.
-On SAEs with an expasion factor of 32 times or more, it can achieve the same or even better performance as the auxiliary loss approach with only half the number of parameters (Section 5.3).
-Broad Architectural Applicability
-The technique maintains applicability to diverse activation functions and architectural variants, as it operates directly on the intrinsic properties of activations.
-This generalizability is further explored in Section 5.4.
+Our approach demonstrates optimal scaling behavior across various SAE training methods. It outperforms TopK and AuxK in any evaluated scale, from 1616K to 11M features (Section 5.3).
+General Applicability
+The technique maintains applicability to diverse architectural variants and activation functions, as it operates directly on the intrinsic properties of activations.
+This generalizability is further explored in Section 5.4 and Appendix H.
 5.3 Scaling Laws
+To understand how our method scales, we evaluate performance as the number of SAE features increases from 1616K to 11M, keeping other hyperparameters fixed (details in Appendix G.5).
 (a) Loss vs. Number of Parameters.
@@
 Details in Section 5.3.
-To understand how our method scales, we evaluate performance as the number of SAE features increases from 1616K to 11M, keeping other hyperparameters fixed (details in Appendix E).
-Figure 7 summarizes the results.
 Active Subspace Init improves reconstruction.
-As shown in Figure 7(a), the normalized MSE follows a smooth power-law decay with increasing feature count.
-Across all scales, Active Subspace Init consistently outperforms baseline TopK and AuxK, achieving superior reconstruction at equivalent feature counts.
+As shown in Figure 7(a), Active Subspace Init consistently outperforms TopK and AuxK across all scales.
 Caveat: some dead features remain at extremely large scales in Active Subspace Init.
@@
 While orthogonal to our initialization method, this choice provides a practical complement that further stabilizes training when scaling SAEs to very large capacities.
-We discuss more about stale momentum and SparseAdam in Appendix A.
+We discuss more about stale momentum and SparseAdam in Appendix C.
 5.4 Generalize to Sparse Replacement Models
@@
 We hypothesize that the low-rank structure of attention outputs contributes significantly to this phenomenon.
-To test this, we apply Active Subspace Initialization to Lorsa [^5] Specifically, we initialize the feature directions (corresponding to $W^{O}$ in Lorsa) within the active subspace.
-Other hyperparameters are the same as He et al.
-(2025)..
-This modification reduces the proportion of dead parameters from $68.4\%$ to $40.5\%$ under identical hyperparameter settings (Table 1), while also slightly improving reconstruction error.
-Table 1: Effect of Active Subspace Initialization on reducing dead parameters in attention replacement model.
-Method Vanilla Active Subspace Init
-Dead Parameters (%) 68.4 40.5
-Normalized MSE 0.130 0.121
-These results indicate that Active Subspace Initialization provides an effective strategy for mitigating dead parameters when training sparse replacement models on low-rank activations.
-We posit that complete elimination of dead heads may require additional mechanisms, such as the tied initialization used in SAEs to ensure alignment between feature encoding and decoding methods666”Match” means encoder can be initialized to predict relatively accurate feature activation values for decoder..
-This approach has been shown to be crucial for reducing dead features in SAEs (Gao et al., 2024), and its absence in Lorsa may limit further improvement.
+To evaluate this, we apply Active Subspace Initialization to Lorsa.
+This modification reduces the proportion of dead parameters significantly under identical hyperparameter settings (Figure 8), while also improving the reconstruction.
+Specifically, we initialize the decoder (corresponding to $W^{O}$ in Lorsa) within the active subspace of each original MHSA head, while the encoder (corresponding to $W^{V}$ in Lorsa) is aligned with the corresponding input-side directions of the decoder.
+Further implementation details are provided in Appendix I.
+Figure 8: Effect of Active Subspace Initialization on reducing dead parameters in attention replacement model.
+6 Discussion
+When to Use Active Subspace Initialization
+For activation sites that do not exhibit clear low-rank structures (e.g., residual stream), our method yields only limited gains (Appendix B.2).
+The singular value spectrum provides a more reliable indicator: when the spectrum decays rapidly in its tail, applying ASI is likely to be effective.
+Causality between Low-Rank Structure and Dead Features
+We observed a strong correlation between low-rank structure and the emergence of dead features (Section 5), but the causal mechanism remains unclear.
+It may be tied to optimization dynamics or feature competition, yet a precise explanation is lacking.
+A deeper understanding of this relationship could also explain why ASI is less effective under high-$\ell_{0}$ warm-up regimes (Appendix H).
+We leave this important line of inquiry to future work.
 7 Conclusion
@@
 The approach generalizes beyond SAEs to sparse replacement models.
-Appendix A Stale Momentum as Another Root Cause of Dead Features
+Appendix A Declaration of LLM Usage
+We acknowledge the use of large language models (LLMs) to assist in the preparation of this manuscript.
+Specifically, ChatGPT and Claude were utilized during the writing and coding process for the exclusive purpose of improving text clarity, grammar, overall fluency, and for generating boilerplate code structures.
+All ideas, theoretical developments, experimental designs, results, analyses, and scientific conclusions are entirely our own.
+The LLMs acted solely as assistive tools and were not involved in any aspect of the intellectual or scientific work.
+Appendix B Ablation Study
+B.1 Active Subspace Init vs Random Subspace Init
+We employ random subspace initialization as a baseline and observe that it consistently degrades SAE training across all metrics, as shown in Figure 9.
+Figure 9: For activations with a full space dimension of 4096, proportion of dead features (left), normalized MSE (mid) and Delta LM loss (right) across different subspace dimensions.
+Random subspaces are used as the baseline, whereas only initialization with the active subspace yields improvement.
+B.2 Apply on Near-Full-Rank Activation
+We also apply Active Subspace Initialization (ASI) to near-full-rank activations, such as those in the residual stream, to evaluate its generality.
+When training an SAE on the post-layer-15 residual stream of Llama-3.1-8B, we find ASI yields minimal gains (Figure 10).
+This is consistent with our expectation, as these activations inherently exhibit a lower rate of dead features even with standard initialization.
+Figure 10: For activations with a full space dimension of 4096, proportion of dead features (left), normalized MSE (mid) and Delta LM loss (right) across different subspace dimensions.
+Random subspaces are used as the baseline, whereas only initialization with the active subspace yields improvement.
+B.3 Choice of the Initial Dictionary Size $D_{init}$
+As shown in Figure 5, $D_{init}$ is a hyperparameter with a wide range of acceptable values (from 256 to 2048).
+We hypothesize that the performance degradation at very low values occurs because an excessively small dictionary fails to span the subspace containing critical information needed for effective reconstruction.
+Appendix C Stale Momentum as Another Root Cause of Dead Features
 Recent work by Bricken et al.
-[2023a] identifies stale momentum as a key cause to dead feature formation.
+(2023a) identifies stale momentum as a key cause to dead feature formation.
 Specifically, when a feature remains inactive over training steps, its associated optimizer momentum continues to accumulate.
@@
 We believe that this is a core technique for scaling sparse dictionary methods, as stale momentum is a common problem for them.
-Appendix B Activation Sources
+Appendix D Activation Sources
 The spectral characteristics of activations vary substantially across model architectures, datasets, and positional contexts.
@@
 Activation Positions
-Unless otherwise specified, activations are extracted from intermediate layers.
+Unless otherwise specified, activations are extracted from model intermediate layers.
 For example, in LLaMA-3.1-8B (32 layers), we use activations from layer 15 (zero-indexed).
-We analyze four types of activations: (1) attention output, (2) MLP output, and (3) residual stream (post layer).
-For each type, we collect 10 million activation vectors.
-We empirically verify that this sample size is sufficient to produce stable and reproducible spectral analyses.
-Appendix C More Low-Rank Result Across Different Models and Datasets
+We analyze three types of activations: (1) attention output, (2) MLP output, and (3) residual stream (post layer).
+Appendix E More Low-Rank Result Across Different Models and Datasets
 (a) Activation spectra for many samples from SlimPajama fed into pythia-2.8b.
@@
 We present relative singular values and fraction of loss recovered for some other model-dataset pairs in Figure 11.
-Models include pythia-2.8b [^10] EleutherAI/pythia-2.8b.
+Models include pythia-2.8b [^11] https://huggingface.co/EleutherAI/pythia-2.8b.
 Datasets include RedPajamaGithub [^12] https://huggingface.co/datasets/cerebras/SlimPajama-627B and CCI3-Data [^13] https://huggingface.co/datasets/BAAI/CCI3-Data
@@
 Figure 12: Comparison of intrinsic dimensions across different variance thresholds.
-Appendix D Different Choose of Variance Threshold for Intrinsic Dimension
+Appendix F Different Choose of Variance Threshold for Intrinsic Dimension
 We use 0.99 as the variance threshold in the main text. We show other threshold chose make no influence to the conclusion in Figure 12. Attention outputs show low-rank structure consistently.
-Appendix E SAE Training Details
-E.1 Collecting Activations
+Appendix G SAE Training Details
+G.1 Collecting Activations
 We truncate each document to 1024 tokens and prepend a <bos> token to the beginning of each document. During training, we exclude the activations corresponding to the <bos> and <eos> tokens.
@@
 Our shuffling strategy maintains a buffer that is reshuffled whenever the buffer is refilled.
-E.2 Initialization and Optimization
-The decoder columns $W^{dec}_{:,i}$ are initialized uniformly and normalized to achieve the lowest initial reconstruction loss.
+G.2 Initialization
+The decoder columns $W^{dec}_{:,i}$ are initialized uniformly, and the optimal norm for them is found through a grid search to minimize the initial reconstruction loss.
 We find that the specific initialization norm has little impact, as long as in a reasonable scope.
 For example, initializing $W^{dec}_{:,i}$ uniformly with a fixed bound, as in Conerly et al.
-[2025], yields similar results.
+(2025), yields similar results.
 The encoder weights $W^{enc}$ are initialized as the transpose of $W^{dec}$, while both the encoder bias $b^{enc}$ and decoder bias $b^{dec}$ are set to zero.
-This initialization scheme ensures that the SAE begins training with an almost zero reconstruction loss. Such initialization has been widely observed to benefit SAE training.
-We train SAEs using the Adam and SparseAdam optimizers, both with $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon=10^{-8}$.
-E.3 Fixed Hyperparameters in Scaling Law
+G.3 Optimization
+We train SAEs using the Adam optimizers, both with $lr=2*10^{-4}$, $batchsize=32768$, $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon=10^{-8}$.
+The differences in scaling law experiments are described in the Appendix G.5.
+G.4 Jumprelu SAEs
+We trained JumpReLU SAEs (Rajamanoharan et al., 2024) under two distinct hyperparameter configurations: one maintaining a consistently low $\ell_{0}$ value throughout training, and another where $\ell_{0}$ is gradually decreased from a higher initial value.
+Unless otherwise specified, all JumpReLU SAEs were trained using the same settings as Conerly et al.
+(2025), which corresponds to the latter configuration.
+The key modifications for the former setting are as follows: (1) we initialized the encoder bias to zero instead of applying the heuristic that equalizes feature activation counts at initialization, and (2) we kept the sparsity coefficient fixed rather than employing a global warm-up schedule.
+As a result, the $\ell_{0}$ sparsity level started at a relatively low value early in training.
+This design is critical to our approach: we observed that if the model remains in a high-$\ell_{0}$ regime (e.g., on the order of $d_{\text{model}}/2$) for an extended period before sparsity increases, the feature directions tend to drift away from the active subspace during this phase, thereby diminishing the effectiveness of our method (Appendix H.1).
+G.5 Fixed Hyperparameters in Scaling Law
 Model, Dataset, Layer, Pos
@@
 Batch Size
-We empirically set $batchsize=4096$, which belows the critical batch size.
+We empirically set the batch size to 40964096—below the critical batch size—which differs from the setting of $32768$ used in other experiments in this paper.
+This reduction is necessary because training SAEs with 1 million features requires substantial memory, posing significant computational challenges.
 Learning Rate
 The learning rate for Adam and SparseAdam is sweeped separately in $1\mathrm{e}{-5}$, $2\mathrm{e}{-5}$, $4\mathrm{e}{-5}$, $6\mathrm{e}{-5}$, $8\mathrm{e}{-5}$, $1\mathrm{e}{-4}$, $2\mathrm{e}{-4}$, $4\mathrm{e}{-4}$], and we ultimately use $4\mathrm{e}{-5}$ for Adam and $6\mathrm{e}{-5}$ for SparseAdam.
+We employ a three-phase learning rate schedule consisting of a linear warm-up, a stable phase, and a linear decay.
+The learning rate increases linearly from zero to its maximum value over the first 500 steps, remains constant during the intermediate phase, and then decays linearly to 1% of the maximum value over the final 20% of the total training steps.
 AuxK
-We follow Gao et al. [2024] to set auxiliary loss coefficient $\alpha$ as $\frac{1}{32}$. We sweep the $k_{aux}$ in [256, 512, 1024, 2048] and finally choose 512.
+We follow Gao et al. (2024) to set auxiliary loss coefficient $\alpha$ as $\frac{1}{32}$. We sweep the $k_{aux}$ in [256, 512, 1024, 2048] and finally choose 512.
 Dimension of Subspace for SAE Initialization
-As shown in Figure 5, $d_{init}$ is a hyperparameter with a wide range of sub-optimal value space (from 256 to 2048). We use 768 for all experiments.
+We use 768 for all experiments.
 Total Tokens
-We use $2.5B$ tokens for each SAE training.
+We use 2.52.5B tokens for each SAE training.
+Appendix H Use ASI on other Activation Functions
+H.1 Jumprelu
+Another wildly used activation fuction is Jumprelu (Rajamanoharan et al., 2024).
+We trained the Jumprelu SAEs under two different hyperparameter settings: one with a consistently low $\ell_{0}$ value and another where $\ell_{0}$ gradually decreases from a higher initial value, as described in Appendix G.4.
+We observed that our method is effective in the former case (Figure 13) but shows little improvement in the latter (Figure 14).
+For cases where one follows a schedule that gradually reduces $\ell_{0}$ from a high initial value, we recommend first applying PCA to reduce the dimensionality of the data.
+The SAE can then be trained on the reduced representation until the $\ell_{0}$ level reaches the target range.
+Afterwards, the PCA projection matrix can be folded into the model parameters, and training can continue in the original space.
+This achieves a similar effect without the drawbacks of prolonged training in the high-$\ell_{0}$ regime.
+Figure 13:
+Figure 14:
+H.2 TopK with K Anneal
+To enhance the finding in Section H.1, we conduct experiments on a variant of TopK, which sets K to a high value and then lets it decrease during training (He et al., 2024).
+We find ASI also fails in this case (Figure 15).
+Figure 15:
+Appendix I Pseudo-code for implementing Active Subspace Init
+Below is a PyTorch-style pseudo-code for Active Subspace Initialization.
+Use on SAE
+# X: activation batch [batch_size, d_model]
+# W_E: decoder weight [d_model, d_sae]
+# W_D: decoder weight [d_sae, d_model], initialized uniformly
+# d_active_subspace: target subspace dimension
+# 1. Demean the activations
+demeaned_X = X - X.mean(dim=0) # [batch_size, d_model]
+# 2. Compute SVD
+U, S, V = torch.svd(demeaned_label) # V: [d_model, d_model]
+# 3. Take top-d_init singular vectors
+proj_weight = V[:, :d_init]  # [d_model, d_init]
+# 4. Fold projection into decoder weights
+W_D.copy_(W_D[:, :d_init] @ proj_weight.T)
+# 5. Init W_E with W_D.T
+W_E.copy_(W_D.T)
+Use on Lorsa
+# X: activation batch [batch_size, seq_len, d_model]
+# mhsa: pretrained MHSA module with W_O and W_V
+# W_O, W_V: Lorsa decoder/encoder weights
+# d_head: dimension of original MHSA heads
+# n_heads: number of original MHSA heads
+# n_lorsa_heads: number of Lorsa heads
+# 1. Run MHSA to capture per-head outputs
+Z = mhsa.compute_z(X) # [batch_size, seq_len, n_heads, d_head]
+output_per_head = torch.einsum(’bsnh, nhd->bsnd’, Z, mhsa.W_O)
+# 2. For each original head
+# initialized some lorsa heads into it’s active subspace
+rate = n_lorsa_heads // n_heads
+for mhsa_index in range(n_heads):
+    head_slice = [rate*mhsa_index:rate*(mhsa_index+1)]
+    # [B, S, d_model]
+    output=output_per_head[:, :, mhsa_index, :]
+    output_flat=output.flatten(0, 1) # [B*S,d_model]
+    # 3.1 Demean
+    demeaned_output=output_flat - output_flat.mean(dim=0)
+    # 3.2 Compute SVD
+    U, S, V = torch.svd(demeaned_output)
+    # 3.3 Take top-d_head singular vectors
+    proj_weight = V[:, :d_head]
+    # 3.4 Update part of decoder weights W_O
+    W_O[head_slice] = W_O[head_slice, :d_head] @ proj_weight.T
+    # 3.5 Update part of encoder weights W_V
+    head_trans = mhsa.W_V[mhsa_index] @ mhsa.W_O[mhsa_index]
+    W_V[head_slice] = W_O[head_slice] @ head_trans.T
+The strategy of initialize $W_{V}$ in Lorsa is a method like the tied initialization used in SAEs to ensure alignment between feature encoding and decoding141414”Match” means encoder can be initialized to predict relatively accurate feature activation values for decoder..
+This approach has been shown to be crucial for reducing dead features in SAEs (Gao et al., 2024).
+We think the same thought could also be used to improve the replacement model for MLP (trancoder and cross layer transcoder), which we leave a deeper investigation to future work.

### v3

@@
-Attention Layers Add Into Low-Dimensional Residual Subspaces
+Dimensional Collapse in Transformer Attention Outputs: A Challenge for Sparse Dictionary Learning
 Abstract
 Transformer architectures, and their attention mechanisms in particular, form the foundation of modern large language models.
-While transformer models are widely believed to operate in high-dimensional hidden spaces, we show that attention outputs are confined to a surprisingly low-dimensional subspace, where about 60% of the directions account for 99% of the variance–a phenomenon that is consistently observed across diverse model families and datasets, and is induced by the attention output projection matrix.
+While transformer models are widely believed to operate in high-dimensional hidden spaces, we show that attention outputs are in fact confined to a surprisingly low-dimensional subspace, with an effective dimensionality of only about $60\%$ of the full space.
+In contrast, MLP outputs and residual streams remain much closer to full-rank, exhibiting effective ranks around $90\%$.
+This striking dimensional discrepancy is consistently observed across diverse model families and datasets, and is strongly shaped by the attention output projection matrix.
 Critically, we find this low-rank structure as a key factor of the prevalent dead feature problem in sparse dictionary learning, where it creates a mismatch between randomly initialized features and the intrinsic geometry of the activation space.
@@
 Our findings provide both new insights into the geometry of attention and practical tools for improving sparse dictionary learning in large language models.
+Machine Learning, ICML
+Figure 1: (left) Attention outputs exhibit pronounced low-rank structure compared to residual streams and MLP outputs, indicating that the attention layer writes into a subspace of the residual stream.
+(right) Low effective dimensionality of activations is a root cause of dead features in sparse dictionary learning methods.
+Setting feature directions in the active subspace mitigates this issue.
 1 Introduction
@@
 As these approaches scale in performance and model coverage, they provide increasingly complete and fine-grained explanations of neural network behavior (Lindsey et al., 2024a; Gao et al., 2024).
-However, scaling these approaches presents practical difficulties (Templeton et al., 2024; Gao et al., 2024; Mudide et al., 2025).
-As models and feature dictionaries grow, the number of parameters increases rapidly, driving up computational costs.
-At the same time, the prevalence of dead features leads to substantial waste in computation and memory (Templeton et al., 2024; Kissane et al., 2024), limiting the efficiency of interpretability methods.
-In this work, we identify low-rank activation structure as a major driver of dead features (Section 5.1).
-In Section 4, we show that attention outputs exhibit a remarkably strong low-rank structure compared to multilayer perceptron (MLP) outputs and residual streams.
-Through singular value decomposition and intrinsic dimensionality analyses (Guth et al., 2023; Staats et al., 2025), we demonstrate that this phenomenon holds universally across layers, datasets, and model families—Llama 3.1 (Dubey et al., 2024), Gemma 2 (Rivière et al., 2024), and Qwen 3 (Yang et al., 2025), which is consistent with the universality hypothesis (Olah et al., 2020; Chughtai et al., 2023; Gurnee et al., 2024; Wang et al., 2025).
-We further trace the origin of this low-rank structure to the anisotropy of the output projection matrix $W^{O}$, which compresses the multi-head outputs into a lower-dimensional subspace.
-In Section 5, we investigate how the low-rank nature of attention outputs interacts with SAE training.
-By evaluating the full suite of open-source SAEs from LlamaScope (He et al., 2024), we show that low intrinsic dimensionality strongly correlates with the number of dead features, suggesting a mismatch between random initialization and the low-dimensional geometry of the activations.
+However, scaling these approaches presents substantial practical challenges (Templeton et al., 2024; Gao et al., 2024; Mudide et al., 2025).
+As models and feature dictionaries grow, the parameter count increases rapidly, leading to significant computational overhead.
+Moreover, a large fraction of learned features remain inactive, resulting in considerable waste in both computation and memory (Templeton et al., 2024; Kissane et al., 2024).
+In this work, we identify the low-rank structure of the activations as a primary driver of dead features (Section 5.1).
+Through singular value decomposition and effective dimensionality analyses (Roy and Vetterli, 2007; Staats et al., 2025), we show for the first time that the outputs of multi-head self-attention in transformer-based language models exhibit a remarkably low-rank structure (Section 4).
+Compared to multilayer perceptron (MLP) outputs and residual streams, attention outputs consistently concentrate in a low-dimensional subspace.
+We show that this behavior is robust across layers, datasets, and model families, including GPT-2 (Radford et al., 2019), Llama 3.1 (Dubey et al., 2024), Gemma 2 (Rivière et al., 2024), and Qwen 3 (Yang et al., 2025).
+This universality aligns with prior observations of shared structures across models (Olah et al., 2020; Chughtai et al., 2023; Gurnee et al., 2024; Wang et al., 2025), while revealing a distinct form of representation collapse (Hua et al., 2021; Jing et al., 2022) at the level of attention outputs.
+We trace the origin of this low-rank structure to the anisotropy of the output projection matrix $W^{O}$, which further compresses multi-head activations into a lower-dimensional subspace.
+In Section 5, we investigate how the low-rank structure of attention outputs interacts with SAE training.
+By evaluating the full suite of open-source SAEs from LlamaScope (He et al., 2024), we show that low effective dimensionality strongly correlates with the number of dead features, suggesting a mismatch between random initialization and the low-dimensional geometry of the activations.
 Drawing inspiration from Phan et al.
@@
 (2024a) and Gao et al.
-(2024), we conduct scaling experiments, which further reveal that ASI achieves superior reconstruction across feature counts, and when combined with SparseAdam [^1] https://docs.pytorch.org/docs/stable/generated/torch.optim.SparseAdam.html, it achieves the best reconstruction in large scale and reduces dead features from 87% to below 1% in Attention Output SAEs with 1M features trained on Llama-3.1-8B (Dubey et al., 2024).
+(2024), we conduct scaling experiments, which further reveal that ASI achieves superior reconstruction across feature counts, and when combined with SparseAdam [^1] See the SparseAdam documentation for details., it achieves the best reconstruction in large scale and reduces dead features from 87% to below 1% in Attention Output SAEs with 1M features trained on Llama-3.1-8B (Dubey et al., 2024).
 Furthermore, we show that Active Subspace Init can generalize to sparse replacement models (He et al., 2025; Dunefsky et al., 2024; Ameisen et al., 2025) (Section 5.4).
@@
 2 Related Work
-2.1 Low-Rankness in Attention Mechanisms
-Prior work has investigated various notions of “low-rankness” within attention mechanisms.: low-rank approximation of attention patterns (Wang et al., 2020; Tay et al., 2020; Raganato et al., 2020), low-rank parameterization for model compression (Noach & Goldberg, 2020; Hu et al., 2022), and the inherent low-rank bottleneck in single-head outputs (Bhojanapalli et al., 2020).
+2.1 Representation Collapse in Neural Representations and Low-Rankness in Attention Mechanism
+A long line of research has shown that neural representations frequently concentrate in low-dimensional subspaces, forming representation collapse (Hua et al., 2021; Tian et al., 2021; Jing et al., 2022).
+These works generally focus on visual models trained using self-supervised methods.
+Within attention mechanisms, prior work has investigated various notions of “low-rankness”: low-rank approximation of attention patterns (Wang et al., 2020; Tay et al., 2020; Raganato et al., 2020), low-rank parameterization for model compression (Noach and Goldberg, 2020; Hu et al., 2022), and the inherent low-rank bottleneck in single-head outputs (Bhojanapalli et al., 2020).
 Different from these prior lines of work, we demonstrate that the multi-head self-attention outputs exhibit a low-rank structure, revealing a distinct and under-explored phenomenon.
-2.2 Linear Representation Hypothesis and Sparse Dictionary Learning Methods
-The linear representation hypothesis suggests that high-level concepts align with linear directions in representation space (Arora et al., 2018; Olah et al., 2020; Elhage et al., 2022; Park et al., 2024).
-Motivated by this view, various sparse dictionary learning methods have been developed for interpretability, including sparse autoencoders and their variants (Cunningham et al., 2023; Bricken et al., 2023b; Lindsey et al., 2024b), transcoders (Dunefsky et al., 2024; Ge et al., 2024), and low-rank sparse attention (He et al., 2025).
-These methods decompose activations into sparse feature combinations, while differing in their strategies for predicting or approximating feature activations.
-The hypothesis has been validated across diverse model scales (Templeton et al., 2024; Lieberum et al., 2024; He et al., 2024), architectures (Wang et al., 2025), and modalities (Abdulaal et al., 2024).
+2.2 Superposition Hypothesis and Sparse Dictionary Learning Methods
+The superposition hypothesis posits that neurons encode multiple non-orthogonal underlying features (Arora et al., 2018; Olah et al., 2020; Elhage et al., 2022; Park et al., 2024).
+Motivated by this view, a variety of sparse dictionary learning methods have been developed for interpretability, including sparse autoencoders and their variants (Cunningham et al., 2023; Bricken et al., 2023b; Lindsey et al., 2024b), transcoders (Dunefsky et al., 2024; Ge et al., 2024), and low-rank sparse attention (He et al., 2025).
+These approaches decompose activations into sparse combinations of learned features while differing in their mechanisms for predicting or approximating feature activations.
+Their successful application across a wide range of model scales (Templeton et al., 2024; Lieberum et al., 2024; He et al., 2024), architectures (Wang et al., 2025), and modalities (Abdulaal et al., 2024) highlights their practical effectiveness for interpretability; however, they do not constitute direct hypothesis tests of the superposition hypothesis, which remain active topics of debate (Sharkey et al., 2025).
 2.3 Dead Features in Sparse Dictionary Learning Methods
@@
 2.4 PCA-Inspired Network Initialization
-A common practice applies PCA to input data for dimensionality reduction before network training (Hastie et al., 2009; Montavon et al., 2012; Jolliffe, 1986; Bishop & Nasrabadi, 2007).
+A common practice applies PCA to input data for dimensionality reduction before network training (Hastie et al., 2009; Montavon et al., 2012; Jolliffe, 1986; Bishop and Nasrabadi, 2007).
 Recently, Phan et al.
@@
 This provides the model with a superior parameter set (Gu et al., 2025), boosting performance by construction.
+Figure 2: Across layers, model families and datasets, attention outputs exhibit dramatically lower effective rank than residual streams and MLP outputs, indicating that the attention layer writing into a low dimensional subspace of residual stream is a universal phenomenon.
+Details in Section 4.1.
+(left) Evaluation of Llama-3.1-8B on SlimPajama (Soboleva et al., 2023) dataset.
+(mid) Middle-layer analysis across model families on SlimPajama dataset.
+(right) Middle-layer analysis of Llama-3.1-8B across datasets.
 3 Preliminaries
@@
 Given input activations $X\in\mathbb{R}^{n\times d}$, where $n$ is the token count and $d$ is the model hidden size, each attention head $i$ computes:
-$Q_{i}=XW^{Q}_{i},\quad K_{i}=XW^{K}_{i},\quad V_{i}=XW^{V}_{i},\quad W^{Q}_{i},W^{K}_{i},W^{V}_{i}\in\mathbb{R}^{d\times d_{h}},$
+$Q_{i}=XW^{Q}_{i}$ $,\qquad K_{i}=XW^{K}_{i},\qquad V_{i}=XW^{V}_{i},$
+$W^{Q}_{i}$ $,\,W^{K}_{i},\,W^{V}_{i}\in\mathbb{R}^{d\times d_{h}},$
 where $d_{h}=d/H$ is the dimensionality of each head, and $H$ is the total number of heads. The attention weights and head outputs are then given by:
-$A_{i}=\mathrm{softmax}\left(\frac{Q_{i}K_{i}^{\top}}{\sqrt{d_{h}}}\right),\quad Z_{i}=A_{i}V_{i}.$
-Let $Z=\text{Concat}[Z_{1},\dots,Z_{H}]\in\mathbb{R}^{n\times d}$ denote the concatenated output of all attention heads (Nanda & Bloom, 2022).
-The final attention output is computed by applying the output projection:
-$O=ZW^{O},\quad W^{O}\in\mathbb{R}^{d\times d}.$
-This formulation shows that $O$ can be viewed as the sum of low-dimensional outputs from each head, projected into the residual stream space.
-$O$ is the attention block’s contribution to the residual stream.
+$\small A_{i}=\mathrm{softmax}\left(\frac{Q_{i}K_{i}^{\top}}{\sqrt{d_{h}}}\right),\quad Z_{i}=A_{i}V_{i}\in\mathbb{R}^{n\times d_{h}}.$
+Let $Z=\text{Concat}[Z_{1},\dots,Z_{H}]\in\mathbb{R}^{n\times d}$ denote the concatenated outputs of all attention heads (Nanda and Bloom, 2022).
+The final attention output is obtained by applying the output projection:
+$O$ $=ZW^{O}=[Z_{1},\dots,Z_{H}]\,[W^{O}_{1},\dots,W^{O}_{H}]^{\top}$
+$=\sum_{i=1}^{H}Z_{i}W^{O}_{i}=\sum_{i=1}^{H}O_{i},$
+where each $W^{O}_{i}\in\mathbb{R}^{d_{h}\times d}$ is the corresponding submatrix of $W^{O}\in\mathbb{R}^{d\times d}$ associated with each head $i$.
+This formulation makes explicit that $O$ is the sum of the outputs from all heads, where each head produces a rank-$d_{h}$ output that is projected into the residual stream space through its corresponding $W^{O}_{h}$.
+Thus, $O$ represents the attention block’s total contribution to the residual stream.
 3.2 TopK Sparse Autoencoders
@@
 Formally, given an input vector $x\in\mathbb{R}^{d}$, the encoder produces
-$z=\mathrm{TopK}(W_{e}x+b_{e}),$
+$\small z=\mathrm{TopK}(W_{e}x+b_{e}),$
 where $\mathrm{TopK}(v)$ sets to zero all but the largest $k$ entries of $v$. The decoder then reconstructs
-$\hat{x}=W_{d}z+b_{d}.$
-The model is trained to minimize the reconstruction loss, optionally augmented with an auxiliary penalty to prevent dead latents:
-$\mathcal{L}_{\text{TopK-SAE}}=\|x-\hat{x}\|_{2}^{2}+\alpha\cdot\mathcal{L}_{\text{aux}},$
+$\small\hat{x}=W_{d}z+b_{d}.$
+The model is trained to minimize the reconstruction loss, optionally augmented with an auxiliary loss to prevent dead latents:
+$\small\mathcal{L}_{\text{TopK-SAE}}=\|x-\hat{x}\|_{2}^{2}+\alpha\cdot\mathcal{L}_{\text{aux}},$
 where $\mathcal{L}_{\text{aux}}$ is an optional term designed to penalize latents that never activate over a training period, and $\alpha$ balances reconstruction fidelity and latent utilization.
 4 Low-Rank Structure of Attention Outputs
-Figure 1: Across layers, model families and datasets, attention outputs exhibit dramatically lower intrinsic dimensions (details in Section 4.1) than residual streams and MLP outputs, showing that the attention layer writing into a low dimensional subspace of residual stream is a universal phenomenon.
 We begin by presenting our central empirical finding: in Transformer models, attention outputs consistently display the strongest low-rank structure compared to MLP outputs and residual streams.
-As shown in Figure 1, attention outputs have a significantly lower intrinsic dimension.
-This phenomenon is remarkably robust, holding across different intermediate layers, model families and datasets.
+As shown in Figure 2, attention outputs have a significantly lower effective rank.
+This phenomenon is remarkably robust, holding across different layers, model families and datasets.
+Further details regarding the activation sources are provided in Section 4.2 and Appendix A.
 These observations highlight that the attention block modifies a subspace of the residual stream, while the MLP operates nearly on the full space.
-4.1 Quantifying Low-Rankness with Relative Singular Values
-We consider activation matrix $\widetilde{A}\in\mathbb{R}^{n\times d}$, where each row corresponds to one token’s activation vector, and $n$ is the number of data points, while $d$ is the dimensionality of the activation space (e.g., the hidden size of the model).
-Unless otherwise specified, $\widetilde{A}$ refers to the mean-centered activations.
-We refer readers to Appendix D for more details of activations sources.
-To quantify the rank of data, we perform singular value decomposition (SVD) on $\widetilde{A}$:
-$\widetilde{A}=U\Sigma V^{\top},$
+4.1 Quantifying Low-Rankness with Effective Rank
+We consider the activation matrix $A\in\mathbb{R}^{n\times d}$, where each row corresponds to the activation vector of a single token.
+Here, $n$ denotes the number of data points and $d$ the dimensionality of the activation space (e.g., the model’s hidden size).
+Unless otherwise specified, $\widetilde{A}$ represents mean-centered activations.
+To quantify the effective dimensionality of the activations, we adopt the effective rank metric introduced by Roy and Vetterli (2007).
+Definition 4.1 (Effective Rank, Roy and Vetterli (2007)).
+Let $\widetilde{A}$ be a nonzero matrix with singular value decomposition $\widetilde{A}=U\Sigma V^{\top}$, where $\Sigma=\mathrm{diag}(\sigma_{1},\sigma_{2},\ldots,\sigma_{r})$ contains singular values in descending order.
+Define the normalized singular value distribution as
+$\small p_{k}=\frac{\sigma_{k}}{\sum_{j=1}^{r}\sigma_{j}},\quad k=1,2,\ldots,r.$
+The (Shannon) entropy of this distribution is
+$\small H(p_{1},p_{2},\ldots,p_{r})=-\sum_{k=1}^{r}p_{k}\log p_{k}.$
+Then, the effective rank of $\widetilde{A}$ is defined as
+$\small\mathrm{erank}(\widetilde{A})=\exp\!\big\{H(p_{1},p_{2},\ldots,p_{r})\big\}.$
+Intuitively, the effective rank captures how evenly the singular values are distributed—higher values indicate a more isotropic spectrum, whereas lower values reflect concentration along a few dominant directions.
+Fraction of effective rank used in Figure 2 means effective rank divided by the dimension of activation space.
+Following Bricken et al.
+(2023b) and Rajamanoharan et al.
+(2024a), we compute the fraction of downstream loss recovered by varying the number of retained components.
+We decompose the activations into singular value components and take the language model loss under full activation ablation as a baseline.
+As components are progressively reintroduced, we report the fraction of this ablated loss that is recovered.
+See Appendix B for a formal definition.
+These quantitative measures complement our core analyses by providing a numerical characterization of the low-rank structure present in activations.
+4.2 Experiment Settings
+For each activation type, we collect a total of 10 million activation vectors.
+[^3] In extremely rare cases, outlier activations inflate variance along certain directions, potentially biasing variance-based dimensionality estimates.
+To mitigate this, we exclude activations whose norms exceed $5\sigma$ from the mean.
+We empirically verified this sample size suffices to ensure stable and reproducible singular spectrum analyses in Appendix C.
+Unless otherwise specified, all experiments in Section 4 run on the middle layer of Llama-3.1-8B (layer 15, zero-indexed), using SlimPajama dataset.
+4.3 Empirical Evidence of Low-Rank Structure
+(a) Singular Value Spectra.
+(b) Fraction of Loss Recovered.
+Figure 3: (a) The attention output is the most low-rank, as indicated by the sharpest decay in singular values. (b) Fraction of loss recovered using varying numbers of top singular components.
+We draw our findings from three lines of evidence:
+Low Effective Rank of Attention Output
+Attention outputs have a effective rank of around 60% of the total dimensionality. In contrast, MLP outputs and the residual streams show much higher effective rank around 90% (Figure 2).
+Rapid Singular Spectral Decay in Attention Output
+This is quantitatively evidenced by the number of components retaining significant energy: only 74.7% singular values exceed 1% [^4] The resolution of bfloat16 (Kalamkar et al., 2019) is 0.01.
+of the maximum in attention output, versus 100.0% for MLP output and residual stream (Figure 3(a)).
+Efficient Downstream Loss Recovery
+Compared to zero ablation, attention output requires only 39.1% of dimensions to recover over 99% of the downstream loss, versus 95.3% and 96.9% of the dimensions for MLP outputs and residual streams to recover the same proportion (Figure 3(b)).
+More results of these metrics across different layers, models, datasets and activation positions are shown in Appendix D.
+4.4 The Output Projection Matrix Further Reduces the Effective Rank of Attention Outputs
+Among all activation types, attention outputs consistently exhibits the most rapid singular spectral decay.
+To investigate whether this low-rank structure originates from the attention heads outputs ($Z$), the output projection matrix ($W^{O}$), or their interaction, we perform a decomposition-based analysis.
+Recall that the attention output is computed as $O=ZW^{O}$, where $Z\in\mathbb{R}^{n\times d}$ is the concatenated output of all attention heads, and $W^{O}\in\mathbb{R}^{d\times d}$ is a learned linear projection.
+To understand how the singular value spectra of $O$ is shaped, we analyze the variance [^5] For zero-mean activations, singular values correspond to the standard deviations of activations along the associated singular directions.
+of $O$ along a unit direction $\hat{e}\in\mathbb{R}^{d}$, given by:
 […]
-To analyze the intrinsic dimension of these activations, we compute the smallest integer $k$ such that:
-$\frac{\sum_{i=1}^{k}\sigma_{i}^{2}}{\sum_{i=1}^{r}\sigma_{i}^{2}}\geq\tau,$
-for a given threshold $\tau\in(0,1)$.
-Since SVD yields the optimal low-rank approximation in terms of reconstruction error, this $k$ provides a principled way to assess how concentrated the activations are in a low-dimensional subspace [^3] We use 0.99 as the threshold in main text, results of some other thresholds are provided in Appendix F, with no influence to the conclusion.
-(Figure 1).
-We further compute the fraction of delta downstream loss recovered by different number of components (Section 4.2).
-These metrics complement our central findings, offering a numerical characterization of low-rankness.
-For each type of activations, we collect 10 million activation vectors [^4] In rare cases, outlier activations inflate variance along some directions, making variance-based dimensionality estimates unreliable; to address this, we filter activations with a norm exceeding $5\sigma$ from the mean..
-We empirically verify that this sample size is sufficient to produce stable and reproducible singular spectrum analysis.
-4.2 Empirical Evidence of Low-Rank Structure
-Figure 2: Results for 1010M tokens from SlimPajama (Soboleva et al., 2023) fed into Llama-3.1-8B.
-(a) The attention outputs are the most low-rank, as indicated by the sharpest decay in singular values.
-(b) Fraction of loss recovered using varying numbers of top singular components.
-We draw our findings from three lines of evidence:
-Low Intrinsic Dimension of Attention Outputs
-Attention outputs has an intrinsic dimension of around 60% of the total dimensionality. In contrast, MLP outputs and the residual streams show much higher intrinsic dimensions above 90% (Figure 1).
-Rapid Singular Spectral Decay in Attention Outputs
-This is quantitatively evidenced by the number of components retaining significant energy: only 74.7% singular values exceed 1% of the maximum in attention outputs, versus 100.0% for MLP outputs and residual streams (Figure 2(a)).
-Efficient Downstream Loss Recovery
-Compared to zero ablation, attention outputs requires only 39.1% of dimensions to recover over 99% of the downstream loss, versus 95.3% and 96.9% for MLP outputs and residual streams (Figure 2(b)).
-More results of these metrics across different layers, models and datasets are shown in Appendix E.
-4.3 Low-Rankness of Attention Outputs Results from the Output Projection Matrix
-Among all activation types, attention outputs consistently exhibits the most rapid spectral decay.
-To investigate whether this low-rank structure originates from the attention heads ($Z$), the output projection matrix ($W^{O}$), or their interaction, we perform a decomposition-based analysis.
-Recall that the attention output is computed as $O=ZW^{O}$, where $Z\in\mathbb{R}^{n\times d}$ is the concatenated output of all attention heads, and $W^{O}\in\mathbb{R}^{d\times d}$ is a learned linear projection.
-To understand how the variance in $O$ (singular value spectra of O) is shaped, we analyze the variance of $O$ along an arbitrary unit direction $\hat{e}\in\mathbb{R}^{d}$, given by:
+This expression highlights that the variance along $e$ is determined by two factors: the norm of $W^{O}\hat{e}$ and the variance of $Z$ along $W^{O}\hat{e}$.
+Specifically, we can rewrite the variance as:
+$\small\mathrm{Var}(O\hat{e})=\mathrm{Var}(Z\hat{v})\cdot\|v\|_{2}^{2}\;\text{where}\;v=W^{O}\hat{e},\ \hat{v}=\frac{v}{\|v\|_{2}}.$
 […]
-This expression highlights that the variance along $e$ is determined by two factors: the norm of $W^{O}e$ and the variance of $Z$ projected onto the direction $W^{O}\hat{e}$.
-Specifically, we can rewrite the variance as:
-$\mathrm{Var}(O\hat{e})=\mathrm{Var}(Z\hat{v})\cdot\|v\|_{2}^{2}\quad\text{where}\quad v=W^{O}\hat{e},\ \hat{v}=\frac{v}{\|v\|_{2}}.$
-We refer to $\mathrm{Var}(Zv)$ as the contribution of Z, capturing how much variance the head output $Z$ provides in that direction, and $\|v\|_{2}^{2}$ as the contribution of $W^{O}$, measuring how much the output projection $W^{O}$ scales or suppresses that direction.
-We compute and visualize both quantities across a set of directions aligned with the right singular vectors of attention output, as shown in Figure 3.
-Our analysis reveals that the low-rank structure of attention outputs $O$ arises primarily from the anisotropy of $W^{O}$, which heavily compresses the output space into a lower-dimensional subspace.
-From a mechanistic perspective, an intuitive way to see this is that although each attention head contributes a $d_{\text{head}}$-dimensional subspace, the superposition of heads (Jermyn et al., 2024; He et al., 2025) inherently leads to overlaps among these subspaces.
-We note the output of the $i^{th}$ head as $\text{head}_{i}$.
+We compute and visualize the singular values of attention output $O$ and these two contributions, as shown in Figure 4.
+This analysis reveals that the low-rank structure of attention outputs is strongly influenced by $W^{O}$, which further compresses $Z$ into a lower-dimensional subspace.
+The analysis of the effective rank of $Z$ in Figure 11 further supports this conclusion.
+From a mechanistic perspective, an intuitive explanation is that although each attention head contributes a $d_{\text{head}}$-dimensional subspace, the superposition of attention heads (Jermyn et al., 2024; He et al., 2025) inevitably induces overlaps among these subspaces.
+Let $O_{i}$ denote the output of the $i^{\text{th}}$ attention head.
 Consequently, the dimension of the MHSA output satisfies
+$\dim\!\left(\bigcup_{i}\text{span}(O_{i})\right)$ $\;\leq\;\sum_{i}\dim\!\big(\text{span}(O_{i})\big)$
+$\;=\;d_{\text{head}}\cdot n_{\text{head}}$
+$\;=\;d_{\text{model}}\qquad\text{(in standard MHSA)}.$
 […]
-$\dim\!\left(\bigcup_{i}\text{span}(\text{head}_{i})\right)\;\leq\;\sum_{i}\dim\!\big(\text{span}(\text{head}_{i})\big)\;=\;d_{\text{head}}\cdot n_{\text{head}}\;(\;=\;d_{\text{model}}\ \text{in standard MHSA)}.$
+Figure 5: The number of dead features (left) and the effective rank (mid) of each activation in Llama-3.1-8B, shows a surprising consistency (right): activations with lower effective rank have more dead features, corresponding to all layers of attention output and last two layers of MLP output.
+Figure 6: After using ASI, proportion of dead features (left), normalized MSE (mid) and Delta LM loss (right) across different $d_{init}$ for activations with a full space dimension of 4096.
+All experiments repeat 3 times using different random seeds and error bars indicate mean ± std.
+Figure 7: At a fixed number of features ($n=32768$), ASI (TopK SAE with Active Subspace Init) achieves a better reconstruction-sparsity trade-off than TopK (standard TopK SAE) and AuxK (TopK SAE with auxiliary loss).
+A similar trend is observed in its impact on Delta LM Loss.
+All experiments repeat 3 times using different random seeds and show the mean.
+The results of std are in Appendix G due to resolution constraints.
 5 Active Subspace Initialization for Sparse Autoencoders
 5.1 Empirical Correlation Between Low-Rank Structure and Dead Features
-Figure 4: The number of dead features (left) and the intrinsic dimension (right) of each activation in Llama-3.1-8B, shows a surprising consistency: activations with lower intrinsic dimensions have more dead features, corresponding to all layers of attention output and last two layers of MLP output.
-To explore how low-rankness affects the interpretability of attention, we use the same framework and data as the original study to evaluate the LlamaScope SAEs (He et al., 2024), which provide a complete set of SAEs trained on attention output, MLP output, and residual stream  [^5] Another prominent open-source SAEs, GemmaScope (Lieberum et al., 2024), train their attention SAEs on Z rather than attention output..
-We find that the number of dead features is strongly related to intrinsic dimensions, as shown in Figure 4.
+To study how low-rankness affects the interpretability of attention, we adopt the same framework and dataset as the original LlamaScope study (He et al., 2024) to evaluate their SAEs trained on attention outputs, MLP outputs, and the residual stream [^6] Another prominent set of open-source SAEs, GemmaScope (Lieberum et al., 2024), train their attention SAEs on Z rather than attention output..
+We find that the number of dead features is strongly related to effective rank, as shown in Figure 5.
 This observation suggests that dead features may stem from the low-rank geometry of the activation space.
+We also train SAEs using different SAE hyperparameters and further systematically verify that this phenomenon is prevalent in Appendix F.
+(a) Loss vs. #Total Features.
+(b) #Dead Features vs. #Total Features
+(c) Loss vs. #Alive Features.
+Figure 8: Scaling results of TopK SAEs and their variants enhanced with AuxK, Active Subspace Init, and SparseAdam–all trained on attention output from the middle layer of Llama-3.1-8B.
+(A) Loss at convergence across different feature counts: Active Subspace Init consistently achieves lower reconstruction error than TopK and AuxK.
+Active Subspace Init with SparseAdam achieves the best at large scale.
+(B) Dead features: Active Subspace Init reduces dead features compared to TopK, but still retains many at extremely large scales.
+Enhanced with SparseAdam, dead features can be reduced to less than 1%.
+(C) Loss across different number of alive features: Active Subspace Init achieves the most efficient utilization of alive features, while AuxK shows the lowest efficiency.
+Details in Section 5.3.
 5.2 Active Subspace Initialization for Sparse Autoencoders
@@
 Let $d$ denote the input dimension, $h$ the hidden dimension of the SAE, and $n$ the number of data points.
-Given activation matrices $\widetilde{A}\in\mathbb{R}^{n\times d}$, we compute the singular value decomposition (SVD):
-$\widetilde{A}=U\Sigma V^{\top},\quad V\in\mathbb{R}^{d\times d},$
-where $U\in\mathbb{R}^{n\times d}$ is the left singular vector matrix, $\Sigma\in\mathbb{R}^{d\times d}$ is the diagonal matrix of singular values, and $V\in\mathbb{R}^{d\times d}$ contains the right singular vectors.
-We select the top $d_{\text{init}}$ right singular vectors to form the active subspace and then initialize the SAE weights directly in this subspace:
-$V_{\text{active}}=V_{:,:d_{\text{init}}}\in\mathbb{R}^{d\times d_{\text{init}}},\quad W_{D}\in\text{span}(V_{\text{active}}),\quad W_{E}=W_{D}^{\top},$
+Given activation matrices $\widetilde{A}\in\mathbb{R}^{n\times d}$ with singular value decomposition $\widetilde{A}=U\Sigma V^{\top}$ and $V\in\mathbb{R}^{d\times d}$ contains the right singular vectors, we select the top $d_{\text{init}}$ singular vectors to define the active subspace:
+$\small V_{\text{active}}=V_{:,:d_{\text{init}}}\in\mathbb{R}^{d\times d_{\text{init}}}.$
+To initialize the SAE within this subspace, we first randomly initialize the decoder weights $W_{D}\in\mathbb{R}^{h\times d}$ and then project their first $d_{\text{init}}$ columns onto the active subspace:
+$\small W_{D}\leftarrow W_{D}\,V_{\text{active}}\,V_{\text{active}}^{\top},\qquad W_{E}=W_{D}^{\top}.$
 where $W_{E}$ is the encoder weight matrix and $W_{D}$ is the decoder weight matrix.
-Intuitively, ASI aligns the initial SAE parameters with the active directions of the data, ensuring that SAEs start in a meaningful low-dimensional subspace.
-As $d_{\text{init}}$ decreases from the full space dimension [^6] Setting $D_{init}$ equal to the full space dimension is equivalent to not using Active Subspace Initialization.
-within a certain range, the number of dead features in the SAE rapidly drops, with a corresponding improvement in Delta LM loss [^7] This metric is defined as the difference between the original language model loss and the loss when the SAE is inserted at the corresponding position, evaluated over 1 million tokens.
+Intuitively, ASI aligns the initial SAE parameters with the effective directions of the data, ensuring that SAEs start in a meaningful low-dimensional subspace.
+As $d_{\text{init}}$ decreases from the full space dimension [^7] Setting $d_{\text{init}}$ equal to the full space dimension is equivalent to not using Active Subspace Initialization.
+within a certain range, the number of dead features in the Attention Output SAE rapidly drops, with a corresponding improvement in Mean Square Error (MSE) and Delta LM loss [^8] Following He et al.
+(2024), this metric is defined as the difference between the original language model loss and the loss when the SAE is inserted at the corresponding position, evaluated over 1 million tokens.
 (Figure 6).
-Further ablation studies and implementation details are contained in Appendix B and Appendix I, respectively.
-Figure 5: After using ASI, proportion of dead features (left), normalized MSE (mid) and Delta LM loss (right) across different subspace dimensions for activations with a full space dimension of 4096.
-Figure 6: At a fixed number of features ($n=32768$), Active Subspace Init achieves a better reconstruction-sparsity trade-off than TopK and AuxK.
-A similar trend is observed in its impact on Delta LM Loss.
-Note: The improvement in downstream loss is less pronounced than that in reconstruction, likely because Active Subspace Init allocates more features to the active subspace, which implicitly enhances reconstruction quality.
+We refer readers to Appendix E for full SAE training details.
+Additional ablations, including the random subspace initialization baseline and the effect of activation rank, are reported in Appendix J.
+Pseudocode for Active Subspace Initialization (ASI) is provided in Appendix K.
 Using Active Subspace Initialization offers several benefits:
@@
 The technique maintains applicability to diverse architectural variants and activation functions, as it operates directly on the intrinsic properties of activations.
-This generalizability is further explored in Section 5.4 and Appendix H.
+This generalizability is further explored in Section 5.4 and Appendix L.
+We conduct a statistical significance test in Appendix H to demonstrate the statistical significance of our conclusions.
+We validate the effectiveness of ASI across different layers, models, and datasets in Appendix I.
+We further compare the features of TopK and ASI in Appendix N, analyzing both the degree of monosemanticity and the behavior of SAE features in the dead subspace, to ensure that ASI increases the number of alive features while preserving feature quality and maintaining the dictionary’s coverage and reconstruction performance in the dead subspace.
+Figure 9: At a fixed number of Lorsa heads ($n=32768$), ASI (TopK Lorsa with Active Subspace Init) achieves a better reconstruction-sparsity trade-off than TopK (standard TopK Lorsa) and AuxK (TopK Lorsa with auxiliary loss).
+A similar trend is observed in its impact on Delta LM Loss.
 5.3 Scaling Laws
-To understand how our method scales, we evaluate performance as the number of SAE features increases from 1616K to 11M, keeping other hyperparameters fixed (details in Appendix G.5).
-(a) Loss vs. Number of Parameters.
-(b) Number of Dead vs. Number of Parameters.
-(c) Loss vs. Number of Alive Parameters.
-Figure 7: Scaling results of TopK SAEs and their variants enhanced with AuxK, Active Subspace Init, and SparseAdam–all trained on attention output from Llama-3.1-8B.
-(A) Loss at convergence across different feature counts: Active Subspace Init consistently achieves lower reconstruction error than TopK and AuxK.
-Active Subspace Init with SparseAdam achieves the best at large scale.
-(B) Dead features: Active Subspace Init reduces dead features compared to TopK, but still retains many at extremely large scales.
-Enhanced with SparseAdam, dead features can be reduced to less than 1%.
-(C) Loss across different number of alive features: Active Subspace Init achieves the most efficient utilization of alive features, while AuxK shows the lowest efficiency.
-Details in Section 5.3.
+To assess scaling, we evaluate our method as the number of SAE features grows from 1616K to 11M, keeping other hyperparameters fixed (see Appendix E).
 Active Subspace Init Improves Reconstruction.
@@
 Building on this insight, we propose using SparseAdam, an optimizer specifically designed for sparse activation settings.
-By updating only the moments and parameters corresponding to non-zero gradients, SparseAdam naturally avoids stale momentum and thus mitigates the dead feature issue.
-As shown in Figures 7(a),  7(b), combining Active Subspace Init with SparseAdam substantially reduces dead features while reaching the lowest reconstruction error.
+By updating only the momentum terms and parameters corresponding to non-zero gradients, SparseAdam avoids stale momentum and thus mitigates the dead feature issue.
+As shown in Figures 8(a),  8(b), combining Active Subspace Init with SparseAdam substantially reduces dead features while reaching the lowest reconstruction error.
 While orthogonal to our initialization method, this choice provides a practical complement that further stabilizes training when scaling SAEs to very large capacities.
-We discuss more about stale momentum and SparseAdam in Appendix C.
-5.4 Generalize to Sparse Replacement Models
+We discuss more about stale momentum and SparseAdam in Appendix M.
+5.4 Generalizing to Sparse Replacement Models
 Recent work by He et al.
-(2025) reports that Lorsa, a sparse replacement model for attention layers, exhibits a high proportion of dead parameters.
-We hypothesize that the low-rank structure of attention outputs contributes significantly to this phenomenon.
-To evaluate this, we apply Active Subspace Initialization to Lorsa.
-This modification reduces the proportion of dead parameters significantly under identical hyperparameter settings (Figure 8), while also improving the reconstruction.
-Specifically, we initialize the decoder (corresponding to $W^{O}$ in Lorsa) within the active subspace of each original MHSA head, while the encoder (corresponding to $W^{V}$ in Lorsa) is aligned with the corresponding input-side directions of the decoder.
-Further implementation details are provided in Appendix I.
-Figure 8: Effect of Active Subspace Initialization on reducing dead parameters in attention replacement model.
-6 Discussion
-When to Use Active Subspace Initialization
-For activation sites that do not exhibit clear low-rank structures (e.g., residual stream), our method yields only limited gains (Appendix B.2).
-The singular value spectrum provides a more reliable indicator: when the spectrum decays rapidly in its tail, applying ASI is likely to be effective.
-Causality between Low-Rank Structure and Dead Features
-We observed a strong correlation between low-rank structure and the emergence of dead features (Section 5), but the causal mechanism remains unclear.
-It may be tied to optimization dynamics or feature competition, yet a precise explanation is lacking.
-A deeper understanding of this relationship could also explain why ASI is less effective under high-$\ell_{0}$ warm-up regimes (Appendix H).
-We leave this important line of inquiry to future work.
+(2025) shows that Lorsa—a sparse replacement for MHSA—has a substantial fraction of dead parameters.
+We hypothesize this is partly due to initialization: standard random initialization ignore the low-dimensional active subspace of attention outputs (Section 4).
+Applying ASI to Lorsa.
+To test whether our subspace-based approach extends beyond SAEs, we incorporate ASI (Section 5.2) into Lorsa training.
+When training Lorsa to approximate a target MHSA, we partition Lorsa heads into such groups, matching the number of attention heads in the MHSA.
+For each group, we initialize the encoder and decoder matrices directly within the corresponding input and output active subspaces of the MHSA head it corresponds.
+In addition, we initialize each group’s Q/K weights from the MHSA Q/K parameters.
+Pseudo code is provided in Appendix K
+Results.
+This initialization sharply reduces dead parameters under identical hyperparameters (Figure 9) while improving reconstruction quality. Further Lorsa training details are in Appendix O.
+6 Discussion and Limitations
+Low-rank attention outputs suggest a new direction for model improvement.
+Modern large language models are predominantly built by stacking attention and MLP blocks with residual connections.
+Our finding that attention outputs show a low-rank structure suggests a potentially new avenue for improving model capacity and expressivity: future architectures may benefit from explicitly mitigating this rank collapse within attention modules.
+Notably, this phenomenon persists across models with and without grouped-query attention (GQA), and under both relative and absolute positional encodings (Section 4), indicating that it may reflect a more fundamental property of the attention mechanism itself rather than an artifact of specific design choices.
+Causality between Low-Rank Structure and Dead Features.
+We find a strong correlation between low-rank activations and the emergence of dead features (Section 5), but the underlying causal mechanism is unresolved.
+This effect may arise from optimization dynamics or feature competition, and we leave a rigorous explanation to future work.
+When to Use Active Subspace Initialization.
+Active Subspace Initialization is most beneficial when activations exhibit pronounced low-rank structure, such as attention outputs, residual streams of some narrow datasets, or certain specialized situations.
+For activation sites without clear low-rank behavior, the improvements are marginal (Appendix J.2).
 7 Conclusion
@@
 The approach generalizes beyond SAEs to sparse replacement models.
-Appendix A Declaration of LLM Usage
-We acknowledge the use of large language models (LLMs) to assist in the preparation of this manuscript.
-Specifically, ChatGPT and Claude were utilized during the writing and coding process for the exclusive purpose of improving text clarity, grammar, overall fluency, and for generating boilerplate code structures.
-All ideas, theoretical developments, experimental designs, results, analyses, and scientific conclusions are entirely our own.
-The LLMs acted solely as assistive tools and were not involved in any aspect of the intellectual or scientific work.
-Appendix B Ablation Study
-B.1 Active Subspace Init vs Random Subspace Init
+Appendix A Activation Sources
+The spectral characteristics of activations vary substantially across model architectures, datasets, and positional contexts.
+Below, we describe the experimental configurations used to support a broad and representative analysis.
+Models
+We study four large language models of different families–GPT-2999https://huggingface.co/openai-community/gpt2, Llama-3.1-8B [^10] https://huggingface.co/meta-llama/Llama-3.1-8B, Qwen3-8B [^11] https://huggingface.co/Qwen/Qwen3-8B, and Gemma-2-9B [^12] https://huggingface.co/google/gemma-2-9b–all based on the Transformer architecture.
+This allows us to assess the robustness of spectral properties under varying model training configurations.
+Datasets
+To investigate how dataset diversity affects activation spectra, we select two datasets with varying linguistic and domain characteristics: (1) SlimPajama [^13] https://huggingface.co/datasets/cerebras/SlimPajama-627B, an English corpus comprising web text, Github, Arxiv and other sources.
+Multi-Corpora in Figure 2 denotes the setting where all SlimPajama components are jointly used with random mixing.
+Github and Arxiv correspond to the code and scientific paper subsets of SlimPajama, respectively.
+(2) CCI3-Data [^14] https://huggingface.co/datasets/BAAI/CCI3-Data, a Chinese dataset with broad domain coverage, which is used in Appendix D as a supplement.
+Activation Positions
+We analyze three types of activations: (1) attention output, (2) MLP output, and (3) residual stream (post layer).
+Appendix B Formal Definition of Fraction of Loss Recovered
+Given the original language model cross-entropy loss is $loss_{\text{original}}$, the loss after ablating the activation at a specific position to zero is $loss_{\text{zero}}$, and the loss after replacing the original activation projected to the subspace spanned by first $n$ singular vectors is $loss_{\text{recovered}}$.
+Then, for these $n$ components, the fraction loss recovered is calculated as:
+$\frac{loss_{\text{zero}}-loss_{\text{recovered}}}{loss_{\text{zero}}-loss_{\text{original}}}$
+Appendix C Error analysis in Singular Value Decomposition
+For the attention output of layer 15 of Llama-3.1-8B, we performed 5 times of singular value decompositions, using different 10 million tokens for each, and calculated the Coefficient of Variation (CV) for each singular value across these 5 runs.
+The maximum CV was only $4.9\times 10^{-3}$, and the mean and standard deviation of the effective rank computed from these 5 SVD results were 2523.165 and 0.404, respectively, with a CV of $1.5\times 10^{-4}$.
+These error experiments show that using 10 million tokens for singular value decomposition is sufficiently stable.
+Appendix D More Singular Spectrum and Effective Rank Results
+(a) Middle layer of pythia-2.8b; SlimPajama
+Effective Dimensionality: Attention Output 1670;
+MLP Output 2252; Residual Stream 2327
+(b) Middle layer of Qwen3-8B; CCI3-Data
+Effective Dimensionality: Attention Output 2356;
+MLP Output 3558; Residual Stream 3140
+(c) Middle layer of Qwen3-8B; SlimPajamaGithub
+Effective Dimensionality: Attention Output 2410;
+MLP Output 3495; Residual Stream 2000
+(d) Middle layer of Qwen3-4B; SlimPajama
+Effective Dimensionality: Attention Output 1122;
+MLP Output 1485; Residual Stream 990
+(e) Middle layer of gpt2; SlimPajama
+Effective Dimensionality: Attention Output 515;
+MLP Output 703; Residual Stream 656
+(f) Middle layer of pythia-160m; SlimPajama
+Effective Dimensionality: Attention Output 434;
+MLP Output 594; Residual Stream 662
+Figure 10:
+D.1 Across Models and Datasets
+We present relative singular values for some other model-dataset pairs in Figure 10.
+Models include pythia-160m [^15] https://huggingface.co/EleutherAI/pythia-160m, pythia-2.8b [^16] https://huggingface.co/EleutherAI/pythia-2.8b.
+Datasets include SlimPajamaGithub (subset of SlimPajama) and CCI3-Data.
+D.2 Across Layers and Activation Positions
+We present the effective rank for activations that are commonly used to train SAEs in Figure 11, including Concatenated Outputs of all Attention Heads (Z), attention output, the hidden activations of MLP (post activation fuction), MLP output, residual stream.
+All effective ranks are computed on SlimPajama.
+(a) The number of the effective rank of each activation in Llama-3.1-8B
+(b) The number of the effective rank of each activation in Gemma-2-9B
+(c) The number of the effective rank of each activation in Qwen3-8B
+Figure 11:
+Appendix E SAE Training Details
+We train SAEs as the following description.
+E.1 Hyperparameters
+Model, Dataset, Layer, Pos
+Llama-3.1-8B, SlimPajama, 15(index start at 0), attention output.
+Sparsity
+We empirically set $k=50$ for a reasonable sparsity following He et al. (2024), except the experiments for sweeping $k$.
+Dictionary Size
+We empirically set $n_{features}=32768$ which is $8\times d_{model}$, except the experiments for sweeping dictionary size (scaling law).
+Batch Size
+We empirically set the batch size to $4096$.
+Optimizer
+We use the Adam and SparseAdam optimizer, both with $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon=10^{-8}$. Unless otherwise specified, Adam is used by default.
+Learning Rate
+The learning rate for Adam and SparseAdam is sweeped separately in $1\mathrm{e}{-5}$, $2\mathrm{e}{-5}$, $4\mathrm{e}{-5}$, $6\mathrm{e}{-5}$, $8\mathrm{e}{-5}$, $1\mathrm{e}{-4}$, $2\mathrm{e}{-4}$, $4\mathrm{e}{-4}$], and we ultimately use $4\mathrm{e}{-5}$ for Adam and $6\mathrm{e}{-5}$ for SparseAdam.
+We employ a three-phase learning rate schedule consisting of a linear warm-up, a stable phase, and a linear decay.
+The learning rate increases linearly from zero to its maximum value over the first 500 steps, remains constant during the intermediate phase, and then decays linearly to 1% of the maximum value over the final 20% of the total training steps.
+AuxK
+We follow Gao et al.
+(2024) to set auxiliary loss coefficient $\alpha$ as $\frac{1}{32}$.
+We sweep the $k_{aux}$ in [256, 512, 1024, 2048] and finally choose 512.
+We also sweep $\alpha$ and find the results are less sensitive to $\alpha$ in a reasonable interval.
+Dimension of Subspace for SAE Initialization ($d_{init}$)
+We use 768 for all experiments, except the experiments for sweeping $d_{init}$. We refer readers to Appendix J.3 for the reason.
+Total Tokens
+We use 800800M tokens for each SAE training.
+E.2 Collecting Activations
+We truncate each document to 1024 tokens and prepend a <bos> token to the beginning of each document. During training, we exclude the activations corresponding to the <bos> , <eos>and <pad> tokens.
+It has been observed that activations from different sequence positions within the same document are often highly correlated and may lack diversity.
+To mitigate this issue, it is common to introduce randomness into the training data.
+Our shuffling strategy maintains a buffer that is reshuffled whenever the buffer is refilled.
+E.3 Initialization
+The decoder columns $W^{dec}_{:,i}$ are initialized uniformly, and the optimal norm for them is found through a grid search to minimize the initial reconstruction loss.
+We find that the specific initialization norm has little impact, as long as in a reasonable scope.
+For example, initializing $W^{dec}_{:,i}$ uniformly with a fixed bound, as in Conerly et al.
+(2025), yields similar results.
+The encoder weights $W^{enc}$ are initialized as the transpose of $W^{dec}$, while both the encoder bias $b^{enc}$ and decoder bias $b^{dec}$ are set to zero.
+E.4 Jumprelu SAEs
+We trained JumpReLU SAEs (Rajamanoharan et al., 2024b) under two distinct hyperparameter configurations: one maintaining a consistently low $\ell_{0}$ value throughout training, and another where $\ell_{0}$ is gradually decreased from a higher initial value.
+Unless otherwise specified, all JumpReLU SAEs were trained using the same settings as Conerly et al.
+(2025), which corresponds to the latter configuration.
+The key modifications for the former setting are as follows: (1) we initialized the encoder bias to zero instead of applying the heuristic that equalizes feature activation counts at initialization, and (2) we kept the sparsity coefficient fixed rather than employing a global warm-up schedule.
+As a result, the $\ell_{0}$ sparsity level started at a relatively low value early in training.
+This design is critical to our approach: we observed that if the model remains in a high-$\ell_{0}$ regime (e.g., on the order of $d_{\text{model}}/2$) for an extended period before sparsity increases, the feature directions tend to drift away from the active subspace during this phase, thereby diminishing the effectiveness of our method (Appendix L.1).
+Appendix F Additional Analysis on Effective Dimensionality and Dead Features
+This section provides additional experimental evidence supporting the claim that low effective dimensionality strongly correlates with a higher proportion of dead features in sparse autoencoders (SAEs).
+In Figure 5, we report the effective rank of the residual stream, attention output, and MLP output at each layer of Llama-3.1-8B, along with the proportion of dead features in the SAEs trained on these activations.
+All SAEs were obtained from Llamascope (He et al., 2024), which uses the same dictionary size (32768 features) and sparsity level ($L_{0}=50$).
+To provide systematic analysis, we additionally train SAEs across multiple dictionary sizes (16384, 32768, 65536) and sparsity levels ($L_{0}\in\{32,64,128\}$) following Appendix E The SAEs are trained on the activations of Llama-3.1-8B, and the corresponding effective ranks of these activations can be found in Figure 4 of the main paper.
+Tables 13 summarize the proportion of dead features across these configurations.
+Across all settings, attention outputs consistently show substantially higher dead-feature ratios than residual streams.
+This trend holds even when the dictionary size varies by a factor of four and the sparsity level varies by a factor of four.
+Table 1: Proportion of dead features for $L_{0}=32$ across different dictionary sizes.
+Activation (Effective Rank) 16384 32768 65536
+Layer 7 attention (2351) 84.80% 90.31% 94.02%
+Layer 7 residual (3664) 1.61% 6.69% 17.10%
+Layer 15 attention (2506) 68.70% 79.86% 87.22%
+Layer 15 residual (3611) 27.01% 45.70% 61.33%
+Layer 23 attention (2654) 58.45% 70.48% 78.81%
+Layer 23 residual (3634) 0.13% 0.26% 1.35%
+Table 2: Proportion of dead features for $L_{0}=64$ across different dictionary sizes.
+Activation (Effective Rank) 16384 32768 65536
+Layer 7 attention (2351) 66.97% 75.65% 82.96%
+Layer 7 residual (3664) 0.02% 0.06% 0.15%
+Layer 15 attention (2506) 41.65% 54.68% 67.43%
+Layer 15 residual (3611) 1.73% 7.96% 18.83%
+Layer 23 attention (2654) 41.58% 56.70% 67.05%
+Layer 23 residual (3634) 0.15% 0.09% 0.08%
+Table 3: Proportion of dead features for $L_{0}=128$ across different dictionary sizes.
+Activation (Effective Rank) 16384 32768 65536
+Layer 7 attention (2351) 49.85% 56.96% 64.83%
+Layer 7 residual (3664) 0.00% 0.00% 0.02%
+Layer 15 attention (2506) 15.09% 25.61% 37.11%
+Layer 15 residual (3611) 0.07% 0.12% 0.44%
+Layer 23 attention (2654) 21.90% 39.02% 52.68%
+Layer 23 residual (3634) 0.14% 0.09% 0.07%
+Appendix G Complete Results of SAE metrics
+We use 3 different random seeds for all experiments in Figure 7 and compute the mean values and the standard deviations of each metrics, as shown in Table 4.
+Table 4: Comparison of Base, AuxK, and ASI across different $L_{0}$ settings. Numbers show mean $\pm$ std over 3 seeds.
+$L_{0}$ Metric Base AuxK ASI
+40 Dead Feature 20395.00 $\pm$ 72.77 37.00 $\pm$ 5.20 10.67 $\pm$ 1.15
+Normalized MSE 0.36323 $\pm$ 0.00018 0.34328 $\pm$ 0.00011 0.33724 $\pm$ 0.00022
+Delta LM Loss ($\times 10^{-3}$) 9.650 $\pm$ 0.040 8.801 $\pm$ 0.053 8.697 $\pm$ 0.033
+50 Dead Feature 16144.33 $\pm$ 129.52 54.67 $\pm$ 5.51 4.00 $\pm$ 1.73
+Normalized MSE 0.33367 $\pm$ 0.00014 0.32241 $\pm$ 0.00020 0.31680 $\pm$ 0.00006
+Delta LM Loss ($\times 10^{-3}$) 8.375 $\pm$ 0.029 7.958 $\pm$ 0.032 7.847 $\pm$ 0.050
+60 Dead Feature 12239.33 $\pm$ 165.51 76.00 $\pm$ 9.17 2.67 $\pm$ 1.53
+Normalized MSE 0.31106 $\pm$ 0.00007 0.30555 $\pm$ 0.00007 0.30000 $\pm$ 0.00005
+Delta LM Loss ($\times 10^{-3}$) 7.503 $\pm$ 0.059 7.325 $\pm$ 0.017 7.214 $\pm$ 0.026
+70 Dead Feature 8854.67 $\pm$ 40.51 115.33 $\pm$ 15.37 1.67 $\pm$ 0.58
+Normalized MSE 0.29295 $\pm$ 0.00008 0.29064 $\pm$ 0.00008 0.28575 $\pm$ 0.00013
+Delta LM Loss ($\times 10^{-3}$) 6.805 $\pm$ 0.074 6.694 $\pm$ 0.021 6.664 $\pm$ 0.066
+80 Dead Feature 6311.33 $\pm$ 55.77 109.67 $\pm$ 11.50 1.00 $\pm$ 1.73
+Normalized MSE 0.27787 $\pm$ 0.00003 0.27715 $\pm$ 0.00003 0.27341 $\pm$ 0.00003
+Delta LM Loss ($\times 10^{-3}$) 6.259 $\pm$ 0.005 6.217 $\pm$ 0.028 6.168 $\pm$ 0.033
+Appendix H Statistical Significance Test
+To assess whether the performance improvements introduced by ASI are statistically significant, we conducted a comprehensive significance analysis across multiple evaluation metrics.
+H.1 Experimental Setup
+We evaluate the statistical significance of performance differences between ASI and two baseline methods (TopK and AuxK) under the following controlled setting:
+  • Model / Layer / $L_{0}$ / Dictionary Size: Llama-3.1-8B, Layer 15, $L_{0}=50$, dictionary size = 32,768.
+  • Number of runs: 15 independent trials for each method, each with a different random seed.
+  • Evaluation metrics: Dead Feature Count, Normalized MSE, and $\Delta$ LM Loss. All metrics follow a “lower is better” criterion.
+  • Comparisons performed: ASI vs. TopK and ASI vs. AuxK for all three metrics.
+H.2 Hypothesis Testing Framework
+For each metric and each baseline method, we perform Welch’s t-test (also known as Welch’s unequal-variance t-test), which does not assume equal variances between groups.
+For ASI and a given baseline method, we test the following hypotheses:
+[…]
+$H_{1}$ $:\mu_{\text{ASI}}<\mu_{\text{baseline}}$ $\text{(ASI outperforms the baseline)}.$
+This is a one-tailed test, as we explicitly test whether ASI achieves significantly lower metric values.
+We use the following SciPy function for all tests:
+$\texttt{scipy.stats.ttest\_ind}(\text{ASI},\text{baseline},\texttt{equal\_var=False},\texttt{alternative='less'}).$
+H.3 Results
+Table 5 reports the resulting p-values for all comparisons. A smaller p-value indicates stronger evidence that ASI outperforms the baseline.
+Table 5: Welch’s t-test results for ASI compared with TopK and AuxK across 15 random seeds.
+Comparison Metric p-value
+ASI vs. TopK Dead Feature Count $3.26\times 10^{-35}$
+Normalized MSE $2.99\times 10^{-40}$
+$\Delta$ LM Loss $1.14\times 10^{-23}$
+ASI vs. AuxK Dead Feature Count $4.33\times 10^{-15}$
+Normalized MSE $6.33\times 10^{-40}$
+$\Delta$ LM Loss $1.27\times 10^{-6}$
+Across all evaluation metrics and both baseline methods, the p-values are far below standard significance thresholds (e.g., $\alpha=0.05$).
+Therefore, we reject the null hypothesis for all comparisons.
+These results demonstrate that the improvements achieved by ASI are statistically significant and robust across random seeds.
+Appendix I Additional Evaluation Across Layers, Models, and Datasets
+To assess the robustness and generality of ASI, we extend our experiments beyond the primary configuration used in the main paper (Llama-3.1-8B, Layer 15, SlimPajama).
+In particular, we investigate whether the advantages of ASI over baseline approaches (TopK and AuxK) persist across different layers, models, and datasets.
+We consider this evaluation essential, as mechanisms in sparse autoencoding can vary substantially across architectural depth, data distribution, and model family.
+I.1 Evaluation on Llama-3.1-8B Across Multiple Layers
+We first evaluate ASI on two additional layers of Llama-3.1-8B (Layers 7 and 23), using the SlimPajama dataset. Activations are taken from the attention output. Results are summarized in Table 6.
+Table 6: Performance of ASI and baseline methods on Llama-3.1-8B Layers 7 and 23. Lower values indicate better performance.
+Dead Feature Count Normalized MSE $\Delta$ LM Loss ($\times 10^{-3}$)
+Layer Base AuxK ASI Base AuxK ASI Base AuxK ASI
+7 25836 98 5308 0.32870 0.29800 0.28882 5.414 4.589 4.430
+23 15542 332 27 0.21302 0.20235 0.20141 1.942 1.865 1.845
+We observe that ASI consistently achieves the lowest reconstruction error (Normalized MSE) and the smallest degradation in language modeling performance ($\Delta$ LM loss).
+For Layer 7, ASI still retains a number of dead features, which may be attributed to its smaller effective rank (2351) compared to Layer 23 (2654).
+Since we use a fixed $d_{\text{init}}$ across layers, this mismatch can lead to remaining dead features.
+Despite this, ASI still achieves the lowest MSE on both layers.
+I.2 Evaluation on Qwen3-8B Across Layers and a New Dataset
+To further test cross-model and cross-dataset generality, we conduct experiments on the Qwen3-8B model using the fineweb-edu dataset, again using attention-output activations.
+Results for Layers 8 and 26 are shown in Table 7.
+Table 7: Performance comparison on Qwen3-8B Layers 8 and 26 using the fineweb-edu dataset. Lower values indicate better performance.
+Dead Feature Count Normalized MSE $\Delta$ LM Loss ($\times 10^{-3}$)
+Layer Base AuxK ASI Base AuxK ASI Base AuxK ASI
+8 19048 4 7 0.31228 0.28849 0.28533 2.9561 2.5911 2.5667
+26 16286 66 566 0.30090 0.28127 0.28087 1.3406 1.2980 1.2922
+The results again confirm that ASI achieves the lowest reconstruction error and the smallest increase in LM loss across both layers.
+The near-dead-feature-free representation produced by AuxK is also observed, but ASI consistently outperforms AuxK in reconstruction quality and LM preservation.
+I.3 Summary
+Across all tested configurations––spanning multiple layers, two large language model families, and two datasets––ASI exhibits consistent advantages over baseline methods.
+These evaluations provide strong empirical evidence that the benefits of ASI are not confined to a specific layer, model, or dataset, but instead generalize across diverse settings.
+Appendix J Ablation Study
+J.1 Active Subspace Init vs Random Subspace Init
 We employ random subspace initialization as a baseline and observe that it consistently degrades SAE training across all metrics, as shown in Figure 12.
@@
 Random subspaces are used as the baseline, whereas only initialization with the active subspace yields improvement.
-B.2 Apply on Near-Full-Rank Activation
+J.2 Apply on Near-Full-Rank Activation
 We also apply Active Subspace Initialization (ASI) to near-full-rank activations, such as those in the residual stream, to evaluate its generality.
@@
 Random subspaces are used as the baseline, whereas only initialization with the active subspace yields improvement.
-B.3 Choice of the Initial Dictionary Size $D_{init}$
+J.3 Choice of the Initial Dictionary Size $d_{init}$
 As shown in Figure 6, $d_{init}$ is a hyperparameter with a wide range of acceptable values (from 256 to 2048).
-We hypothesize that the performance degradation at very low values occurs because an excessively small dictionary fails to span the subspace containing critical information needed for effective reconstruction.
-Appendix C Stale Momentum as Another Root Cause of Dead Features
-Recent work by Bricken et al.
-(2023a) identifies stale momentum as a key cause to dead feature formation.
-Specifically, when a feature remains inactive over training steps, its associated optimizer momentum continues to accumulate.
-If the feature activates, the stale momentum results in disproportionately large updates, destabilizing training and potentially suppressing that feature permanently.
-To directly address this, we adopt SparseAdam, an optimizer tailored for sparse activation settings, designed for more efficient use of compute and memory.
-SparseAdam updates both parameters and moments only when the corresponding feature is active.
-This could effectively prevent the harmful accumulation of stale momentum.
-Empirically, we observe that this change substantially reduces the rate of dead feature formation in large-scale SAE training.
-We believe that this is a core technique for scaling sparse dictionary methods, as stale momentum is a common problem for them.
-Appendix D Activation Sources
-The spectral characteristics of activations vary substantially across model architectures, datasets, and positional contexts.
-Below, we describe the experimental configurations used to support a broad and representative analysis.
-Models
-We study three large language models of different families–Llama-3.1-8B [^8] https://huggingface.co/meta-llama/Llama-3.1-8B, Qwen3-8B [^9] https://huggingface.co/Qwen/Qwen3-8B, and Gemma-2-9B [^10] https://huggingface.co/google/gemma-2-9b–all based on the Transformer architecture.
-This allows us to assess the robustness of spectral properties under varying model training configurations.
-Datasets
-To investigate how dataset diversity affects activation spectra, we select three datasets with varying linguistic and domain characteristics: (1) SlimPajama, an English corpus comprising web text, books, and other sources; (2) RedPajamaGithub, a large-scale code corpus; and (3) CCI3-Data, a Chinese dataset with broad domain coverage.
-Activation Positions
-Unless otherwise specified, activations are extracted from model intermediate layers.
-For example, in LLaMA-3.1-8B (32 layers), we use activations from layer 15 (zero-indexed).
-We analyze three types of activations: (1) attention output, (2) MLP output, and (3) residual stream (post layer).
-Appendix E More Low-Rank Result Across Different Models and Datasets
-(a) Activation spectra for many samples from SlimPajama fed into pythia-2.8b.
-(b) Activation spectra for many samples from CCI3-Data fed into Qwen3-8B.
-(c) Activation spectra for many samples from RedPajamaGithub fed into Qwen3-8B.
-Figure 11:
-We present relative singular values and fraction of loss recovered for some other model-dataset pairs in Figure 11.
-Models include pythia-2.8b [^11] https://huggingface.co/EleutherAI/pythia-2.8b.
-Datasets include RedPajamaGithub [^12] https://huggingface.co/datasets/cerebras/SlimPajama-627B and CCI3-Data [^13] https://huggingface.co/datasets/BAAI/CCI3-Data
-(a) Intrinsic dimension with threshold $0.5$.
-(b) Intrinsic dimension with threshold $0.9$.
-(c) Intrinsic dimension with threshold $0.999$.
-(d) Intrinsic dimension with threshold $0.9999$.
-Figure 12: Comparison of intrinsic dimensions across different variance thresholds.
-Appendix F Different Choose of Variance Threshold for Intrinsic Dimension
-We use 0.99 as the variance threshold in the main text. We show other threshold chose make no influence to the conclusion in Figure 12. Attention outputs show low-rank structure consistently.
-Appendix G SAE Training Details
-G.1 Collecting Activations
-We truncate each document to 1024 tokens and prepend a <bos> token to the beginning of each document. During training, we exclude the activations corresponding to the <bos> and <eos> tokens.
-It has been observed that activations from different sequence positions within the same document are often highly correlated and may lack diversity.
-To mitigate this issue, it is common to introduce randomness into the training data.
-Our shuffling strategy maintains a buffer that is reshuffled whenever the buffer is refilled.
-G.2 Initialization
-The decoder columns $W^{dec}_{:,i}$ are initialized uniformly, and the optimal norm for them is found through a grid search to minimize the initial reconstruction loss.
-We find that the specific initialization norm has little impact, as long as in a reasonable scope.
-For example, initializing $W^{dec}_{:,i}$ uniformly with a fixed bound, as in Conerly et al.
-(2025), yields similar results.
-The encoder weights $W^{enc}$ are initialized as the transpose of $W^{dec}$, while both the encoder bias $b^{enc}$ and decoder bias $b^{dec}$ are set to zero.
-G.3 Optimization
-We train SAEs using the Adam optimizers, both with $lr=2*10^{-4}$, $batchsize=32768$, $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon=10^{-8}$.
-The differences in scaling law experiments are described in the Appendix G.5.
-G.4 Jumprelu SAEs
-We trained JumpReLU SAEs (Rajamanoharan et al., 2024) under two distinct hyperparameter configurations: one maintaining a consistently low $\ell_{0}$ value throughout training, and another where $\ell_{0}$ is gradually decreased from a higher initial value.
-Unless otherwise specified, all JumpReLU SAEs were trained using the same settings as Conerly et al.
-(2025), which corresponds to the latter configuration.
-The key modifications for the former setting are as follows: (1) we initialized the encoder bias to zero instead of applying the heuristic that equalizes feature activation counts at initialization, and (2) we kept the sparsity coefficient fixed rather than employing a global warm-up schedule.
-As a result, the $\ell_{0}$ sparsity level started at a relatively low value early in training.
-This design is critical to our approach: we observed that if the model remains in a high-$\ell_{0}$ regime (e.g., on the order of $d_{\text{model}}/2$) for an extended period before sparsity increases, the feature directions tend to drift away from the active subspace during this phase, thereby diminishing the effectiveness of our method (Appendix H.1).
-G.5 Fixed Hyperparameters in Scaling Law
-Model, Dataset, Layer, Pos
-Llama-3.1-8B, SlimPajama, 15(index start at 0), attention output.
-Sparsity
-We empirically set $k=50$ for a reasonable sparsity in scaling laws.
-Batch Size
-We empirically set the batch size to 40964096—below the critical batch size—which differs from the setting of $32768$ used in other experiments in this paper.
-This reduction is necessary because training SAEs with 1 million features requires substantial memory, posing significant computational challenges.
-Learning Rate
-The learning rate for Adam and SparseAdam is sweeped separately in $1\mathrm{e}{-5}$, $2\mathrm{e}{-5}$, $4\mathrm{e}{-5}$, $6\mathrm{e}{-5}$, $8\mathrm{e}{-5}$, $1\mathrm{e}{-4}$, $2\mathrm{e}{-4}$, $4\mathrm{e}{-4}$], and we ultimately use $4\mathrm{e}{-5}$ for Adam and $6\mathrm{e}{-5}$ for SparseAdam.
-We employ a three-phase learning rate schedule consisting of a linear warm-up, a stable phase, and a linear decay.
-The learning rate increases linearly from zero to its maximum value over the first 500 steps, remains constant during the intermediate phase, and then decays linearly to 1% of the maximum value over the final 20% of the total training steps.
-AuxK
-We follow Gao et al. (2024) to set auxiliary loss coefficient $\alpha$ as $\frac{1}{32}$. We sweep the $k_{aux}$ in [256, 512, 1024, 2048] and finally choose 512.
-Dimension of Subspace for SAE Initialization
-We use 768 for all experiments.
-Total Tokens
-We use 2.52.5B tokens for each SAE training.
-Appendix H Use ASI on other Activation Functions
-H.1 Jumprelu
-Another wildly used activation fuction is Jumprelu (Rajamanoharan et al., 2024).
-We trained the Jumprelu SAEs under two different hyperparameter settings: one with a consistently low $\ell_{0}$ value and another where $\ell_{0}$ gradually decreases from a higher initial value, as described in Appendix G.4.
-We observed that our method is effective in the former case (Figure 13) but shows little improvement in the latter (Figure 14).
-For cases where one follows a schedule that gradually reduces $\ell_{0}$ from a high initial value, we recommend first applying PCA to reduce the dimensionality of the data.
-The SAE can then be trained on the reduced representation until the $\ell_{0}$ level reaches the target range.
-Afterwards, the PCA projection matrix can be folded into the model parameters, and training can continue in the original space.
-This achieves a similar effect without the drawbacks of prolonged training in the high-$\ell_{0}$ regime.
-Figure 13:
-Figure 14:
-H.2 TopK with K Anneal
-To enhance the finding in Section H.1, we conduct experiments on a variant of TopK, which sets K to a high value and then lets it decrease during training (He et al., 2024).
-We find ASI also fails in this case (Figure 15).
-Figure 15:
-Appendix I Pseudo-code for implementing Active Subspace Init
+We hypothesize that the performance degradation when $d_{init}$ is very low is due to a combination of the dictionary failing to cover the subspace containing the key information needed for effective reconstruction and the features being too crowded.
+Appendix K Pseudo-code for implementing Active Subspace Init
 Below is a PyTorch-style pseudo-code for Active Subspace Initialization.
@@
 # 4. Fold projection into decoder weights
-W_D.copy_(W_D[:, :d_init] @ proj_weight.T)
+W_D.copy_(W_D @ proj_weight @ proj_weight.T)
 # 5. Init W_E with W_D.T
@@
 Use on Lorsa
-# X: activation batch [batch_size, seq_len, d_model]
-# mhsa: pretrained MHSA module with W_O and W_V
-# W_O, W_V: Lorsa decoder/encoder weights
-# d_head: dimension of original MHSA heads
-# n_heads: number of original MHSA heads
-# n_lorsa_heads: number of Lorsa heads
-# 1. Run MHSA to capture per-head outputs
-Z = mhsa.compute_z(X) # [batch_size, seq_len, n_heads, d_head]
-output_per_head = torch.einsum(’bsnh, nhd->bsnd’, Z, mhsa.W_O)
-# 2. For each original head
-# initialized some lorsa heads into it’s active subspace
-rate = n_lorsa_heads // n_heads
-for mhsa_index in range(n_heads):
-    head_slice = [rate*mhsa_index:rate*(mhsa_index+1)]
-    # [B, S, d_model]
-    output=output_per_head[:, :, mhsa_index, :]
-    output_flat=output.flatten(0, 1) # [B*S,d_model]
-    # 3.1 Demean
-    demeaned_output=output_flat - output_flat.mean(dim=0)
-    # 3.2 Compute SVD
-    U, S, V = torch.svd(demeaned_output)
-    # 3.3 Take top-d_head singular vectors
-    proj_weight = V[:, :d_head]
-    # 3.4 Update part of decoder weights W_O
-    W_O[head_slice] = W_O[head_slice, :d_head] @ proj_weight.T
-    # 3.5 Update part of encoder weights W_V
-    head_trans = mhsa.W_V[mhsa_index] @ mhsa.W_O[mhsa_index]
-    W_V[head_slice] = W_O[head_slice] @ head_trans.T
-The strategy of initialize $W_{V}$ in Lorsa is a method like the tied initialization used in SAEs to ensure alignment between feature encoding and decoding141414”Match” means encoder can be initialized to predict relatively accurate feature activation values for decoder..
+# Input:
+# X: input activations [b, s, d]  (b=batch_size, s=seq_len, d=d_model)
+# mhsa: pretrained MHSA module
+#   mhsa.W_V: [n, d, h]  (n=n_heads, h=d_head, note: n*h=d)
+#   mhsa.W_O: [n, h, d]
+# Lorsa parameters to initialize:
+#   W_V: [n_lorsa, d]  (n_lorsa = number of Lorsa heads)
+#   W_O: [n_lorsa, d]
+# d_qk: Lorsa head dimension for initialization
+# 1. Compute per-head V projections
+X_flat = X.reshape(b*s, d)                         # [b*s, d]
+W_V_cat = mhsa.W_V.permute(1,0,2).reshape(d, d)    # [d, d]
+V_per_head = (X_flat @ W_V_cat).reshape(b*s, n, h) # [b*s, n, h]
+# 2. Project V back to d_model space for each head
+# captured_v[:, i, :] = V_per_head[:, i, :] @ mhsa.W_V[i].T
+captured_v = einsum(’bnh, nhd -> bnd’,
+                    V_per_head, mhsa.W_V.permute(0,2,1))
+# captured_v: [b*s, n, d]
+# 3. Initialize Lorsa heads from each original head’s active subspace
+rate = n_lorsa // n
+for i in range(n):
+    slice_i = [rate*i : rate*(i+1)]
+    # 3.1 Extract this head’s captured V
+    v = captured_v[:, i, :]            # [b*s, d]
+    # 3.2 Demean
+    demeaned_v = v - v.mean(dim=0)     # [b*s, d]
+    # 3.3 SVD on transposed data to get principal directions
+    U, S, _ = svd(demeaned_v.T)        # demeaned_v.T: [d, b*s]
+                                       # U: [d, d]
+    # 3.4 Take top-d_qk principal directions as projection
+    proj = U[:, :d_qk]                 # [d, d_qk]
+    # 3.5 Update W_V: project from initial d_qk space to principal subspace
+    W_V[slice_i] = W_V[slice_i, :d_qk] @ proj.T    # [rate, d]
+    # 3.6 Update W_O: chain updated W_V through original head’s OV circuit
+    # OV_i = mhsa.W_V[i] @ mhsa.W_O[i]: [d, h] @ [h, d] = [d, d]
+    W_O[slice_i] = W_V[slice_i] @ mhsa.W_V[i] @ mhsa.W_O[i]
+    # [rate, d] @ [d, h] @ [h, d] = [rate, d]
+# 4. Normalize all Lorsa weights (row-wise)
+W_V = W_V / W_V.norm(dim=1, keepdim=True)   # [n_lorsa, d]
+W_O = W_O / W_O.norm(dim=1, keepdim=True)   # [n_lorsa, d]
+The strategy of initialize $W_{O}$ in Lorsa is a method like the tied initialization used in SAEs to ensure alignment between feature encoding and decoding171717”Match” means encoder can be initialized to predict relatively accurate feature activation values for decoder..
 This approach has been shown to be crucial for reducing dead features in SAEs (Gao et al., 2024).
 We think the same thought could also be used to improve the replacement model for MLP (trancoder and cross layer transcoder), which we leave a deeper investigation to future work.
+Appendix L Use ASI on other Activation Functions
+L.1 Jumprelu
+Another wildly used activation fuction is Jumprelu (Rajamanoharan et al., 2024b).
+We trained the Jumprelu SAEs under two different hyperparameter settings: one with a consistently low $\ell_{0}$ value and another where $\ell_{0}$ gradually decreases from a higher initial value, as described in Appendix E.4.
+We observed that our method is effective in the former case (Figure 14) but shows little improvement in the latter (Figure 15).
+We train these SAEs following Appendix E.
+For cases where one follows a schedule that gradually reduces $\ell_{0}$ from a high initial value, we recommend first applying PCA to reduce the dimensionality of the data.
+The SAE can then be trained on the reduced representation until the $\ell_{0}$ level reaches the target range.
+Afterwards, the PCA projection matrix can be folded into the model parameters, and training can continue in the original space.
+This achieves a similar effect without the drawbacks of prolonged training in the high-$\ell_{0}$ regime.
+Figure 14: For the attention output of the middler layer of Llama-3.1-8B, using ASI on JumpRelu SAEs which has a low initial $\ell_{0}$ is effective. Details in Appendix L.1
+Figure 15: For the attention output of the middler layer of Llama-3.1-8B, using ASI on JumpRelu SAEs which has a high initial $\ell_{0}$ than gradually decreasing shows little improvement.
+Details in Appendix L.1
+L.2 TopK with K Anneal
+To enhance the finding in Section L.1, we conduct experiments on a variant of TopK, which sets K to a high value and then lets it decrease during training (He et al., 2024).
+We find ASI also fails in this case (Figure 16).
+Figure 16: For the attention output of the middler layer of Llama-3.1-8B, using ASI on TopK SAEs which sets K to a high value and then lets it decrease during training fails. Details in Appendix L.2
+Appendix M Stale Momentum as Another Root Cause of Dead Features
+Recent work by Bricken et al.
+(2023a) identifies stale momentum as a key cause to dead feature formation.
+Specifically, when a feature remains inactive over training steps, its associated optimizer momentum continues to accumulate.
+If the feature activates, the stale momentum results in disproportionately large updates, destabilizing training and potentially suppressing that feature permanently.
+To directly address this, we adopt SparseAdam, an optimizer tailored for sparse activation settings, designed for more efficient use of compute and memory.
+SparseAdam updates both parameters and moments only when the corresponding feature is active.
+This could effectively prevent the harmful accumulation of stale momentum.
+Empirically, we observe that this change substantially reduces the rate of dead feature formation in large-scale SAE training.
+We believe that this is a core technique for scaling sparse dictionary methods, as stale momentum is a common problem for them.
+Appendix N Compare Features of SAEs with and without ASI
+N.1 Monosemanticity
+We conducted an additional analysis to assess the degree of monosemanticity exhibited by features learned by the base TopK SAE and the ASI-enhanced SAE.
+Following the rubric of Cunningham and Conerly (2024), we performed a blinded evaluation of 100 randomly sampled features and recorded the semantic consistency scores assigned to each feature.
+For clarity, we reproduce below the scoring rubric used for evaluating activation consistency:
+  • 5: Clear pattern with no deviating examples
+  • 4: Clear pattern with one or two deviating examples
+  • 3: Clear overall pattern but quite a few examples not fitting that pattern
+  • 2: Broad consistent theme but lacking structure
+  • 1: No discernible pattern
+To provide transparency, Tables 8 summarize the distribution of scores for TopK and ASI.
+Across both models, we found no statistically significant differences in score distributions at this scale of analysis.
+Variations between the two SAE variants were marginal and did not indicate systematic differences in feature quality.
+This result is consistent with our expectations, as our method does not modify the SAE architecture and is not designed to intervene in how features are formed.
+Consequently, the potential risk of degrading feature quality remains low.
+Table 8: Monosemanticity score distributions for ASI-enhanced SAE and base TopK SAE features.
+(a) ASI-enhanced SAE
+Score Count
+5 17
+4 14
+3 6
+2 5
+1 8
+(b) Base TopK SAE
+Score Count
+5 18
+4 12
+3 6
+2 6
+1 8
+N.2 Analysis of SAE Features in the Dead Subspace
+To assess whether ASI alters the SAE’s behavior in directions corresponding to the dead subspace, we perform a comparative analysis between Attention Output SAEs trained with and without ASI under identical configurations (same number of features, $K$, and all other hyperparameters).
+Feature alignment with the dead subspace.
+We compute the cosine similarity between each SAE feature and the dead subspace, restricting the analysis to alive features (which accounts for the difference in total counts).
+The distribution of cosine values is summarized in Table 9.
+Across all intervals, the ASI-initialized SAE exhibits a larger number of alive features, while both methods exhibit very few features that align closely with the dead subspace.
+This suggests two possible explanations: (i) features in the dead subspace have extremely small magnitude and provide insufficient signal for the SAE to learn, or (ii) the dead subspace does not contain meaningful standalone features, and only small components of features reside in this region.
+Table 9: Distribution of cosine similarity between SAE features and the dead subspace (alive features only).
+Method [0.0, 0.05) [0.05, 0.1) [0.1, 0.15) [0.15, 1]
+TopK 10176 189 6 0
+ASI 24576 936 8 0
+Reconstruction error in the dead subspace.
+We further project the reconstruction error onto the dead subspace to quantify the SAE’s reconstruction performance on components lying in this region.
+As shown in Table 10, the reconstruction errors are nearly identical between the two methods, with the ASI showing only a marginal improvement.
+Table 10: Reconstruction error projected onto the dead subspace.
+Method MSE in dead subspace
+TopK 0.00350
+ASI 0.00334
+Overall, these analyses indicate that ASI has minimal impact on the SAE’s behavior within the dead subspace, while substantially reducing the number of dead features.
+Appendix O Lorsa Implementation Details
+O.1 Hyperparameters
+Model, Dataset, Layer
+Llama-3.1-8B, SlimPajama, 15(index start at 0).
+Dictionary Size
+We empirically set the num of Lorsa heads $n_{heads}^{Lorsa}=32768$ which is $8\times d_{model}$.
+We set the num of QK group of Lorsa $n_{qk}^{Lorsa}=256$ which is $8\times n_{heads}^{MHSA}$.
+We set the dimension of QK of Lorsa $d_{qk}^{Lorsa}=128$ which is $d_{qk}^{MHSA}$.
+Batch Size
+We empirically set the batch size to $32768$.
+Optimizer
+We use the Adam optimizer, with $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon=10^{-8}$.
+Learning Rate
+The learning rate is sweeped in $1\mathrm{e}{-5}$, $2\mathrm{e}{-5}$, $4\mathrm{e}{-5}$, $6\mathrm{e}{-5}$, $8\mathrm{e}{-5}$, $1\mathrm{e}{-4}$, $2\mathrm{e}{-4}$, $4\mathrm{e}{-4}$], and we ultimately use $2\mathrm{e}{-4}$.
+We employ a three-phase learning rate schedule consisting of a linear warm-up, a stable phase, and a linear decay.
+The learning rate increases linearly from zero to its maximum value over the first 500 steps, remains constant during the intermediate phase, and then decays linearly to 1% of the maximum value over the final 20% of the total training steps.
+AuxK
+We follow Gao et al. (2024) to set auxiliary loss coefficient $\alpha$ as $\frac{1}{32}$. We sweep the $k_{aux}$ in [256, 512, 1024, 2048] and finally choose 512.
+Dimension of Subspace for SAE Initialization ($d_{init}$)
+Because the active subspace of the input and output of each MHSA heads is very close to the dimension of MHSA head ($d_{head}$), we set it directly to $d_{head}$.
+We found that increasing or decreasing this value did not improve performance.
+Total Tokens
+We use 800800M tokens for each Lorsa training.
+Sequence Length
+We truncate each document to 2048 tokens. During training, we exclude the activations corresponding to the <bos> , <eos>and <pad> tokens.
+O.2 Initialization
+We initialize the query and key matrices $W_{Q}$ and $W_{K}$ using Xavier uniform initialization (Glorot and Bengio, 2010).
+The value matrix $W_{V}$ is initialized from a normal distribution $\mathcal{N}(0,1/\sqrt{d_{sae}})$, while the output matrix $W_{O}$ is initialized from $\mathcal{N}(0,1/\sqrt{d_{model}})$.
+All bias terms $b_{Q}$, $b_{K}$, $b_{V}$, and $b_{D}$ (if used) are initialized to zero.